### v1

# UltraWiki: Ultra-fine-grained Entity Set Expansion with Negative Seed Entities

Abstract.
Entity Set Expansion (ESE) aims to identify new entities belonging to the same semantic class as a given set of seed entities. Traditional methods primarily relied on positive seed entities to represent a target semantic class, which poses challenge for the representation of ultra-fine-grained semantic classes. Ultra-fine-grained semantic classes are defined based on fine-grained semantic classes with more specific attribute constraints. Describing it with positive seed entities alone cause two issues: (i) Ambiguity among ultra-fine-grained semantic classes. (ii) Inability to define “unwanted” semantic. Due to these inherent shortcomings, previous methods struggle to address the ultra-fine-grained ESE (Ultra-ESE). To solve this issue, we first introduce negative seed entities in the inputs, which belong to the same fine-grained semantic class as the positive seed entities but differ in certain attributes. Negative seed entities eliminate the semantic ambiguity by contrast between positive and negative attributes. Meanwhile, it provide a straightforward way to express “unwanted”. To assess model performance in Ultra-ESE and facilitate further research, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. UltraWiki encompasses 50,973 entities and 394,097 sentences, alongside 236 ultra-fine-grained semantic classes, where each query of them is represented with 3-5 positive and negative seed entities. A retrieval-based framework RetExpan and a generation-based framework GenExpan are proposed to comprehensively assess the efficacy of large language models from two different paradigms in Ultra-ESE. Moreover, we devised three strategies to enhance models’ comprehension of ultra-fine-grained entities semantics: contrastive learning, retrieval augmentation, and chain-of-thought reasoning. Extensive experiments confirm the effectiveness of our proposed strategies and also reveal that there remains a large space for improvement in Ultra-ESE.
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/THUKElab/UltraWiki.
1. Introduction
Entity Set Expansion (ESE) is a critical task aiming to identify new entities belonging to the same semantic class as a given set of seed entities (Zhang et al., 2020; Li et al., 2022a, 2023d, 2023g). ESE plays an important role for a wide range of user-tailored applications (Li et al., 2023b; Yu et al., 2023; Li et al., 2022c; Zhang et al., 2023b), such as recommendation systems (Jacucci et al., 2021; Ma et al., 2022; Ye et al., 2023), knowledge base population (Zhu et al., 2023; Velardi et al., 2013; Li et al., 2023a, f, e) and domain text analysis (Kohita et al., 2020; Liu et al., 2022; Li et al., 2023c, 2022b), thus demanding a high degree of semantic class granularity.
Traditional methods of ESE primarily relied on positive seed entities to represent a target semantic class. However, solely leveraging positive seed entities poses a significant challenge in representing ultra-fine-grained semantic classes. Fine-grained semantic classes are concrete at the concept level, e.g., mobile phone brands, car brands, and the target entities of different semantic classes are typically quite different. In contrast, ultra-fine-grained semantic classes are defined based on concepts with more specific attribute constraints, such as mobile phone brands using Android, Asian mobile phone brands. Notably, ultra-fine-grained semantic classes within the same fine-grained class often demonstrate significant overlap in target entities. As shown in Figure 1, describing the target semantic class merely with positive seed entities cause two issues: (i) Ambiguity among ultra-fine-grained semantic classes. For instance, distinguishing between Mobile phone brands using Android and American Mobile phone brands using only positive seeds like {Motorola, Microsoft Mobile, Google} may lead to ambiguous expansion results. These seeds do not capture the distinct features that delineate each class. (ii) Inability to define “unwanted”. Positive seeds alone cannot represent the “unwanted” semantic. Considering when users wish to find Mobile phone brands not made by Asian manufacturers, it is impractical to define this semantic class by enumerating phone brands from other continents (Europe, North America, Africa, etc.) in the positive seeds. Since the size of the positive seeds is limited, and it’s also unreasonable to expect users to specify all potential attribute values.
Figure 1. Semantic classes at different levels of granularities, and issues with positive seed entity only.
Previous exploratory works mainly made efforts towards fine-grained expansion. ProbExpan (Li et al., 2022a) refines fine-grained semantic boundaries by mining in-batch hard negative samples. FGExpan (Xiao et al., 2023) guides the expansion process to yield more precise results for the fine-grained semantic class with the existing taxonomy. MultiExpan (Li et al., 2023g) leverages the visual information of entities to differentiate between semantic classes. Nonetheless, these methods primarily concentrate on conceptual level distinctions (Although these concepts are somehow hard to distinguish, e.g., US states vs. US cities). They are insufficient for distinguishing ultra-fine-grained semantic classes, due to the inherent shortcoming of representing semantic classes with positive seeds only.
Inspired by exploiting the negative feedback to improve the modeling of user purchase intent in the recommendation system (Zhao et al., 2018; Wu et al., 2020; Xie et al., 2021), we solve ultra-fine-grained ESE (Ultra-ESE) by introducing negative seed entities in the inputs. Specifically, the input negative seed entity belongs to the same fine-grained semantic class as the positive seed entity but differs in certain attributes. Positive and negative seed entities jointly represent and constrain target semantic classes, which addresses the representation granularity challenge head-on: Firstly, the contrast between positive and negative seeds on attributes can express the attribute of user interests, eliminating the semantic class ambiguity caused by positive seeds alone. For instance, for the aforementioned example, when the negative seeds are entities of Mobile phone brands using iOS, it indicates the user’s concern about the operation system, irrespective of the manufacturer. Secondly, negative seeds provide a straightforward way to express “unwanted” semantics.
Regrettably, current ESE dataset (Shen et al., 2017; Pradhan et al., 2013) lacks the concept of negative seed entities, let alone ultra-fine-grained semantic classes. To bridge this gap, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. Derived from Wikipedia, UltraWiki encompasses 10 fine-grained semantic classes, alongside 50,973 entities and 394,097 sentences. We determined 2-3 attributes for each fine-grained semantic class and annotated each entity with attribute values utilizing a semi-automatic approach. Based on the identified attributes, we further constructed 236 ultra-fine-grained semantic classes, in which each semantic class is represented with 3-5 positive and negative seed entities, respectively. In other words, there is an average of 23 (236/10-1) hard negative semantic classes for each ultra-fine-grained semantic class, which poses a significant challenge to the model’s capability for ultra-fine-grained semantic comprehension. Even advancing GPT-4 can’t handle it well.
In experiments, we evaluate the performance of existing methods on UltraWiki. Furthermore, we designed a retrieval-based framework with Encoder-only models (BERT (Devlin et al., 2018)) and a generation-based framework with Decoder-only models (LLaMA (Touvron et al., 2023)) to comprehensively assess the efficacy of large language models (LLMs) from two different paradigms on the Ultra-ESE task. Meanwhile, we proposed three strategies to enhance models’ ability to comprehend ultra-fine-grained semantics of entities: contrastive learning, retrieval augmentation, and chain-of-thought reasoning.
In summary, the main contributions include:
  • We proposed the more challenging ultra-fine-grained ESE task, and introduce negative seed entities to represent ultra-fine-grained semantic classes more precisely for the first tiem.
  • We constructed the first large-scale dataset UltraWiki, tailored for ultra-fine-grained ESE tasks. It encompasses 10 fine-grained semantic classes and 261 ultra-fine-grained semantic classes.
  • We designed both retrieval-based and generation-based frameworks to assess the efficacy of BERT-like and GPT-like LLMs on the UltraWiki dataset. Furthermore, three strategies are proposed for refining the semantic comprehension of ultra-fine-grained entities.
  • Extensive experiments confirm the effectiveness of our proposed strategies and also reveal significant potential for enhancing ultra-fine-grained semantic comprehension of entities by LLMs.
2. Related Work
2.1. Methods of ESE
Recently, increasing emphasis has been placed on the expansion of fine-grained (e.g., US city) rather than coarse-grained (e.g., Location) semantic classes. ProbExpan (Li et al., 2022a) employs heuristic methods to mine hard negative entities of target semantic classes, thereby refining the semantic boundaries of fine-grained classes through contrastive learning (Chen et al., 2020; Gao et al., 2021). FGExpan (Xiao et al., 2023) leverages the existing taxonomy to direct BERT in reasoning about more fine-grained names of semantic classes. Besides, MultiExpan (Li et al., 2023g) integrates multimodal pre-trained models to encode visual information (images) associated with entities, which benefits the differentiation of various fine-grained semantic classes. Nevertheless, these methods do not essentially tackle the issue that solely using positive seed entities fails to sufficiently represent ultra-fine-grained semantic classes. In contrast, our research introduces the novel concept of incorporating negative seed entities alongside positive ones to describe an ultra-fine-grained semantic class.
It is worth mentioning that there has been some work to incorporate negative entities (Curran et al., 2007; Jindal and Roth, 2011; Shi et al., 2014; Li et al., 2022a), though the usage of negative entities in these methods is relatively naive. We point out that the role of the negative entities used in previous work is fundamentally different from ours. Negative entities in previous work are purely used to help determine the boundary of the target set described by positive seeds. In contrast, negative entities in our model are used to describe the target ultra-fine-grained classes that cannot be characterized by positive seeds alone.
2.2. Data Resources of ESE
Queries in the existing ESE dataset merely consist of positive seed entities. The approach of representing ultra-fine-grained semantic classes solely via positive seed entities inherently faces two major issues: (i) Ambiguity among ultra-fine-grained semantic classes. (ii) Inability to define “unwanted” semantics. Consequently, existing datasets lack ultra-fine-grained semantic classes. For instance, Wiki (Shen et al., 2017) and APR (Shen et al., 2017), derived from the Wikipedia and Reuters corpora respectively, encompass merely 3 and 8 fine-grained semantic classes. Some named entity recognition datasets are directly used as ESE evaluation benchmarks, they contain semantic classes with coarser granularity. For example, the OntoNotes (Pradhan et al., 2013) and CoNLL (Sang and De Meulder, 2003) datasets include semantic classes like organization and location. In contrast, UltraWiki encompasses 261 ultra-fine-grained semantic classes, which surpasses existing ESE datasets in term of both granularity and quantity of semantic classes.
Figure 2. llustration of the UltraWiki dataset construction process.
Figure 3. Distribution of positive and negative target entities in two different cases.
3. Task Formulation
Our study focuses on the ultra-fine-grained ESE task, whose input incorporates three components: query $S$, candidate entity vocabulary $V$, and corpus $D$. The query $S=\{S^{pos}\cup S^{neg}\}$ is a composite set comprising positive seed entities $S^{pos}=\{e^{pos}_{1},\ldots,e^{pos}_{k}\}$ and negative seed entities $S^{neg}=\{e^{neg}_{1},\ldots,e^{neg}_{k}\}$, which jointly represent a target ultra-fine-grained semantic class. Both $S^{pos}$ and $S^{neg}$ belong to the same fine-grained semantic class. The entities in $S^{pos}$ and $S^{neg}$ share same values for certain attributes $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$, respectively. Notably, as shown in Figure3, $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$ are not limited in quantity and may overlap or differ. The vocabulary $V$ enumerates the candidate entities. Additionally, the corpus $D$ supplies contextual information $\{e_{i},s^{i}_{1},\ldots,s^{i}_{n}\}$ for each entity $e_{i}$, in which $s^{i}_{n}$ is a sentence containing entity $e_{i}$.
The ultimate objective of ultra-fine-grained ESE is to identify positive target entities from the vocabulary $V$ that belong to the same fine-grained semantic class, and share the same attribute values with $S^{pos}$ while distinct from $S^{neg}$. In an ideal entity feature space, entities that share more same attribute values shall be positioned closer.
4. UltraWiki Dataset
Considering the lack of negative seed entities and ultra-fine-grained semantic classes in existing ESE datasets, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. This section outlines the construction process of UltraWiki. Semantic granularity, annotation consistency, and candidate entities’ difficulty are fully considered to guarantee the quality of UltraWiki.
4.1. Dataset Construction
Two methods exist for constructing UltraWiki. A more direct method is to first collect a large number of entities and corresponding context sentences. Then, for each fine-grained semantic class, the human annotator traverses the entire entity vocabulary to identify the corresponding entity and annotates the attribute values, thereby constructing the ultra-fine-grained semantic class. However, despite the availability of numerous entity-centric datasets, this method is labor-intensive and prone to omissions of entities. Consequently, we chose a more practical strategy. We begin by identifying fine-grained semantic classes along with their corresponding entities, then collecting corresponding sentences and attribute information for each entity in turn. Figure 2 illustrates the entire process.
Step 1. Semantic Classes and Entities Collection. Wikipedia maintains an extensive catalog of entities associated with fine-grained semantic classes. We carefully chose 10 semantic classes with certain principles (discussed in Section 3.2) and then crawled the corresponding entities. Additionally, a substantial number of entities sampled from Wikipedia pages were incorporated into the candidate entity vocabulary as negative entities.
Step 2. Entity-Labeled Sentences Collection. A significant volume of text is also crawled from Wikipedia pages, where entities are uniquely identified by hyperlinks. Since the entities obtained in Step 1 also contain hyperlinks, we can readily establish associations between entities and the sentences containing them, thereby providing abundant contextual information about the entities.
Step 3. Entity Attribute Annotation. For each fine-grained semantic class $C$, we manually select $k$ independent attributes $\mathcal{A}=\{a_{1},a_{2},\cdots,a_{k}\}$, which will be used to generate ultra-fine-grained semantic class in next step. Attributes of each class are shown in appendix. Subjective attributes are avoided like good or bad. We initially employed a script to automatically query Wikidata for the corresponding attribute value. For the remaining attributes that could not be automatically annotated, we employed professional annotators to manually label them, ensuring each entity was reviewed three times for accuracy and consistency.
Step 4. Negative-aware Semantic Classes Generation. We devised an algorithm for automated generation of ultra-fine-grained semantic classes based on positive and negative attributes. Specifically, for attributes $\mathcal{A}$ of each fine-grained semantic class $C$, we sample $n^{pos}$ and $n^{neg}$ attributes as positive and negative attributes (denoted as $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$). Subsequently, we pick a value for all the attributes in $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$. This process yields a list of positive values $\mathcal{V}^{pos}$ and a list of negative values $\mathcal{V}^{neg}$, which jointly constrain the attributes of the target semantic class to achieve ultra-fine-grained control of the semantic class. When $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$ are the same, the role of $\mathcal{A}^{neg}$ is to emphasize attributes concerned by user and mitigate ambiguity. Conversely, when they differ, $\mathcal{A}^{neg}$ serves to convey unwanted semantics. We denote the set of entities satisfying $\mathcal{V}^{pos}$ and $\mathcal{V}^{neg}$ as positive target entity $\mathcal{P}$ and negative target entity $\mathcal{N}$ respectively, ensuring that both $|\mathcal{P}|$ and $|\mathcal{N}|$ exceed the minimum entity requirement $n_{thred}=6$.
Figure 4. Heat map visualization of semantic class similarity.
4.2. Dataset Analysis
Statistics of UltraWiki. UltraWiki is the first large-scale dataset to feature ultra-fine-grained semantic classes. It comprises a collection of 50,973 entities and 394,097 sentences sourced from Wikipedia. UltraWiki encompasses 10 fine-grained semantic classes that comprehensively cover five major entity types, including Organization, Location, Product, Person, and Miscellaneous. Leveraging the combination of positive and negative attributes, we further derived 261 ultra-fine-grained semantic classes. On average, each ultra-fine-grained semantic class comprises 63 positive target entities ($\mathcal{P}$) and 60 negative target entities ($\mathcal{N}$) that are not expected to be expanded. Each ultra-fine-grained semantic class encompasses 3 queries, with 3 to 5 positive seed entities and 3 to 5 negative seed entities. The majority of the semantic classes are constrained by one positive and one negative attribute.
Table 1. Comparison of ESE datasets.
Wiki APR CoNLL ONs UltraWiki
# Semantic Classes 8 3 4 8 261
Semantic granularity Fine Fine Coarse Coarse Ultra-Fine
# Queries per Class 5 5 1 1 3
# Pos Seeds per Query 3 3 10 10 3-5
# Neg Seeds per Query N/A N/A N/A N/A 3-5
# Candidate Entities 33K 76K 6K 20K 51K
# Sentences of Corpus 973K 1043K 21K 144K 394K
Entity Attribution
Quality of UltraWiki. The fine-grained semantic classes, entity corpus, and partial attributes of UltraWiki are automatically crawled from Wikipedia and Wikidata. Both sources are high-quality knowledge bases, curated by numerous domain experts, thereby significantly guaranteeing the quality of UltraWiki. For attributes that require manual annotation, we ensure that each value is annotated by three annotators. Eventually, the inter-annotator agreement measured by Fleiss’s Kappa (Fleiss, 1971) reaches 0.90, indicating satisfying consistency and accuracy of the annotation.
Difficulty of UltraWiki. Ultra-ESE is an inherently challenging task. The challenge arises from the substantial entity overlap among ultra-fine-grained semantic classes ( approximately 99% of ultra-fine-grained semantic classes intersect), which requires the model to comprehend the ultra-fine-grained distinctions of entities across varied contexts. Even advanced models like GPT-4 struggle to address this challenge (refer to the results in the Experimental Section). Additionally, our selected semantic classes includes a subset of long-tail entities, such as lesser-known Chinese cities, which typically lack extensive contextual knowledge. Simultaneously, employing the BM25-based search, we incorporated entities highly similar to the target entities as hard negative entities in the candidate entity vocabulary. The inclusion of these entity types contributes to the difficulty of UltraWiki. The visualization of entity similarity in Figure 4 illustrates that each semantic class of UltraWiki possesses extremely high intra-class similarity.
5. Methods
This section presents our proposed retrieval-based and generation-based frameworks to leverage both positive and negative seed entities, namely RetExpan and GenExpan. Here, we re-emphasize that the inputs of the two frameworks comprise both positive and negative seed entities, which belong to the same fine-grained semantic class and solely differ in their ultra-fine-grained attributes.
Most prior ESE models are built upon retrieval-based frameworks, e.g., CGExpan and ProbExpan. The retrieval-based framework first encodes entities as feature vectors in a low-dimensional semantic space, and then measures the probability that a candidate entity belongs to the target semantic class via the entity representations similarity, which is naturally more suitable for retrieval tasks like ESE. Considering the recent emergence of generative LLMs, we also propose a generation-based framework to explore the potential of generative LLMs. Moreover, for each framework, two separate enhancement strategies are crafted to enhance the models’ capacity for ultra-fine-grained semantic comprehension of entities. The proposed RetExpan and GenExpan provide comprehensive and strong baselines for Ultra-ESE.
Figure 5. Overall Frameworks of RetExpan and GenExpan.
5.1. Retrieval-based Framework: RetExpan
5.1.1. Overall Framework
RetExpan is structured into three steps: entity representation, entity expansion, and entity re-ranking. In the first stage, we design the entity encoder that extracts contextual feature vectors of entities in sentences. An entity is represented as the mean of the feature vectors derived from all sentences containing it. Entity prediction task is introduced to refine the entity representation. During the second stage, RetExpan obtains a preliminary list of target entities based on their similarity to the positive seed entities. Notably, negative seed entities are omitted in this process to ensure the recall of all entities satisfying fine-grained semantic classes. In the third stage, negative seed entities are also employed to re-rank the entity list obtained in the previous phase, reducing the ranking of entities aligning with the attributes of the negative seed entities.
Entity Representation. The goal of entity encoder is to extract contextual features of entities from textual data. Initially, we replace entity mentions within a sentence with [MASK] tokens to construct the input for the encoder. Given a contextual sentence $T$ with masked entity mentions, we utilize the $\text{BERT}_{\text{BASE}}$, which comprises a 12-layer Transformer (Vaswani et al., 2017), to extract contextual embeddings:
(1) $H=\{{h}_{1},{h}_{2},...,{h}_{L}\}=\text{BERT}_{\text{BASE}}(T)$
where $L$ is the length of the tokenized sentences. Ultimately, the contextual feature $\mathbf{h}$ of an entity is computed as the average of the contextual embedding mask $h_{[\text{MASK}]}$ at the mask position across all sentences containing it.
Inspired by ProbExpan, entity prediction task is introduced to refine the entity representation. Concretely, a classification header $\mathbf{f}$ is appended to the entity encoder. After obtaining the hidden state $h_{[\text{MASK}]}$ at [MASK] position, it is transformed into the probability distribution of the masked entity among the candidate entities via MLP and SoftMax:
(2) $\small\hat{y}=\mathbf{f}\left(h_{[\text{MASK}]}\right)=\operatorname{Softmax}\left(\operatorname{MLP}\left(h_{[\text{MASK}]}\right)\right),\hat{y}\in\mathbb{R}^{V_{e}}$
where $V_{e}$ is the size of candidate entities. Then, cross-entropy loss with label smoothing (Szegedy et al., 2016) is applied to learn the latent semantics of entities:
(3) $\small\begin{split}\mathcal{L}_{mask}=-\frac{1}{N}\sum_{i}^{N}\sum_{j}^{V_{e}}y_{i}[j]\cdot(1-\eta)\cdot log(\hat{y}_{i}[j])\\ +(1-y_{i}[j])\cdot\eta\cdot log(1-\hat{y}_{i}[j])\end{split}$
in which ground-truth $y$ is a one-hot vector and $N$ is the batch size. Smoothing factor $\eta$ mitigates over-penalization for entities that exhibit similar semantic to the ground-truth entity.
Entity Expansion. We acquire the preliminary entity list comprising entities that belong to the same fine-grained semantic class as the positive seed entities ${S}^{pos}$. For each candidate entity $e$, we define its positive similarity score $sco^{pos}$ as follows:
(4) $\small sco^{pos}(e)=\frac{1}{|{S}^{pos}|}\sum_{e^{\prime}\in{S}^{pos}}\mathrm{cos}\_\mathrm{sim}\left(h(e),h(e^{\prime})\right)$
We keep top-$K$ entities with the highest $sco^{pos}$ as the initial expansion list $L_{0}$.
Entity Re-ranking. This step further considers negative seed entities, excluding entities that share semantics with them (i.e., possess the same value on negative attribute). Analogous to Equation 4, we define the negative similarity score ${sco}^{neg}$ between candidate entity and negative seed entities. However, directly re-ranking $L_{0}$ in ascending order utilizing $sco^{neg}$ introduces a significant number of noisy entities that do not belong to the same fine-grained semantic class as the seed entities. These irrelevant entities, characterized by low similarity scores with the negative seed entities, may be erroneously assigned over-high ranking.
To address this issue, we propose a simple yet effective strategy: segmented re-ranking. This approach divides $L_{0}$ equally into $\left\lceil\frac{\left|L_{0}\right|}{l}\right\rceil$ segments and then conducts a descending re-ranking using $sco^{neg}$ for each segment individually. Hyper-parameter $l$ represents the length of each segment. This strategy facilitates a local fine ranking of $L_{0}$ and prevents the assignment of over-high rankings to noisy entities with quite low $sco^{neg}$ values.
5.1.2. Strategy 1: Ultra-fine-grained Contrastive Learning
To enhance the ultra-fine-grained semantic comprehension of entities, we devised attribute-aware contrastive learning. Conventional contrastive learning facilitates clearer semantic boundaries for fine-grained semantic classes by drawing the representation of the same semantic class entities closer and the representation of different semantic class entities further apart. Although proven to be effective, basic contrastive learning is not precise enough for Ultra-ESE as it neglects entity differences inside the same semantic class. For instance, ”Samsung” and ”Motorola” may serve as positive pairs in terms of the attribute ”Operating System” but as negative pairs for the attribute ”Manufacturer”. Ideally, entities with more attributes in common should exhibit closer distance in the entity space. To this end, we construct ultra-fine-grained training data based on positive and negative seed entities to implicitly incorporate attribute-based differences into entity representations.
Ultra-fine-grained Training Data. We prompt GPT-4111All specific prompts used for querying GPT-4 can be found in the Appendix. We use GPT-4-1106-preview. to deduce potential positive (negative) attributes and return T entities from the initial list $L_{0}$ that are most similar to the given positive (negative) seed entities to form the entity list $L_{pos}$ ($L_{neg}$). Entities in $L_{pos}$ and $L_{neg}$ are roughly considered to belong to the positive (target) and negative ultra-fine-grained semantic class $\mathcal{P}$ and $\mathcal{N}$, respectively. Hence, to distinguish entities with different attribute values under the same fine-grained semantic class, entities in $L_{pos}$ and $L_{neg}$ should be pulled away from each other. Specifically, we construct the contrastive training pairs as follows:
(5) $\small\begin{split}\mathbb{P}_{pos}=\{(x,x^{\prime})|e(x)\in L_{pos},e(x^{\prime})\in L_{pos}\}\cup\\ \{(x,x^{\prime})|e(x)\in L_{neg},e(x^{\prime})\in L_{neg}\}\cup\{(x,x^{\prime})|e(x)=e(x^{\prime})\},\end{split}$
(6) $\small\begin{split}\mathbb{P}_{neg}=\{(x,y)|e(x)\in L_{pos},e(y)\in L_{neg}\}\cup\\ \{(x,z)|e(x)\in L_{pos}\cup L_{neg},e(z)\in\overline{L_{0}}\},\end{split}$
where $e(x)$ indicates the entity of the training sample $x$. $\overline{L_{0}}$ denotes entities from other fine-grained semantic classes, which form normal negative pairs with entities from $L_{pos}$ and $L_{neg}$. $\overline{L_{0}}$ is necessary to ensure that from the fine-grained semantic level, entities in $L_{pos}$ and $L_{neg}$ remain similar in the semantic space. Moreover, during the training process, we append the corresponding positive and negative seed entities after each training sample (i.e., sentence) to implicitly specify the corresponding ultra-fine-grained semantics, avoiding positive-negative conflicts for the same entity pair.
Contrastive Loss. We choose the classical InfoNCE loss (Oord et al., 2018) as the contrastive loss. Contrastive learning is conducted in a new hypersphere space to prevent semantic collapse, which is transformed by another MLP-based mapping head $\mathbf{f_{cl}}$ and $l$-2 normalization.
5.1.3. Strategy 2: Entity-based Retrieval Augmentation
The contextual information of entities is sourced from the provided corpus ${D}$, which might not offer adequate information for ultra-fine-grained semantic distinctions, particularly for long-tailed entities. Hence, we gather the corresponding introduction text of entities from the external corpus, Wikidata, to augment the entity knowledge. During both training and inference, these entity introductions serve as prefixes for all sentences, thereby enriching the model’s ultra-fine-grained comprehension of the entities.
5.2. Generation-based Framework: GenExpan
5.2.1. Overall Framework
To explore the potential of emerging generative LLMs, we devised the generative framework GenExpan. Unlike RetExpan which relies on intermediate semantic features of entities, GenExpan directly performs inference based on the given preceding text and the knowledge stored in the LLMs themselves, thus is more efficient. The proposed GenExpan consists of three phases: entity generation, entity selection, and entity re-ranking. In the first phase, GenExpan generates a new set of entities based on positive seed entities and current expanded entities. A constrained decoding strategy is employed to ensure that the new entities are included in the candidate entities. Subsequently, the entity selection phase filters the top-$K$ entities to be added to the current expansion, based on the similarity scores between the new entities and the positive seed entities. These first two steps are executed iteratively for several rounds. Lastly, akin to RetExpan, the final step re-ranks the previous expansion results to lower the ranking of entities identical to the negative seed entities in terms of negative attributes. GenExpan utilizes LLaMA-7B as the backbone. Additionally, we continually pre-trained LLaMA-7B to fully leverage the given corpus ${D}$.
Figure 6. Prefix tree-based constrained decoding.
Entity Generation. We craft a simple prompt $\text{Prompt}_{\text{g}}$ with the given entity set to guide the model toward generating entities semantically similar to the target semantic class. The full prompt can be found in appendix. In the first round of generation, we randomly select 3 entities from positive seed entities $S^{pos}$ as the input entity set. In subsequent rounds, to maintain diversity while ensuring the semantic does not deviate from original positive seed entities, 2 entities are randomly chosen from $S^{pos}$, and 1 entity is randomly sampled from current expansion results.
Vanilla decoding strategies like beam search (Sutskever et al., 2014) might generate entities that are not included in the candidate entities. Hence, we construct a prefix tree structure to ensure that the decoded entities belong to the candidate entities. In this structure, the root node represents the beginning, and each path from the root to a leaf node represents a complete candidate entity. During decoding, the process must follow a specific path from root to leaf at a time. In other words, for a certain node, its child nodes represent subsequent tokens that are allowed to be generated. By employing this prefix-constrained beam search, we ensure that the generated entities remain valid.
Entity Selection. For each previously generated new entity $e$, we compute the positive similarity score $sco^{pos}$ between them and positive seed entities $S^{pos}$, which is calculated as follows:
(7) $\small sco^{pos}(e)=\frac{1}{|S^{pos}|}\sum_{e^{\prime}\in S^{pos}}\sqrt[|e^{\prime}|]{P(e^{\prime}|f(e))},$
in which $f(e)$ is the template “{e} is similar to”. $P(e^{\prime}|f(e))$ represents the conditional probability of LLM generating $e^{\prime}$ when given $f(e)$. We use geometric mean to balance the different token numbers $|e^{\prime}|$ of various entities. Top-$p$ entities with the highest positive similarity scores are incorporated into the current expansion result.
Entity Re-ranking. The process of reranking in GenExpan is identical to that in RetExpan except for the calculation of $sco^{neg}$.
5.2.2. Strategy 1: Chain-of-thought Reasoning
Inspired by (Wei et al., 2022), we enhance the reasoning capability of LLMs for Ultra-ESE through a series of intermediate reasoning steps. Basically, the positive seed entities $S^{pos}$ in GenExpan are directly utilized to construct $\text{Prompt}_{\text{g}}$ for generating new entities. However, in chain-of-thought reasoning, we prompt the LLM to initially generate potential fine-grained class names of the positive seed entities, along with the positive attributes that share the same value. This information is then integrated into $\text{Prompt}_{\text{g}}$ to guide subsequent entity generation. The chain-of-thought reasoning stimulates the reasoning capability of LLM and benefits it for perceiving the semantic classes of seed entities and the concerned attributes.
5.2.3. Strategy 2: Entity-based Retrieval Augmentation
Similar to RetExpan, we prepend the introduction text of the input entity before $\text{Prompt}_{\text{g}}$ to enrich the LLM’s understanding of the input entity. Importantly, these entity introductions are not integrated during the fine-tuning of LLaMA; they are exclusively utilized during entity generation.
6. Experiments
6.1. Experiment Setup
Compared Methods. We compare with three categories of models. The initial category encompasses traditional statistical probability-based methods, notably SetExpan (Shen et al., 2017) and CasE (Yu et al., 2019). The second category comprises methods based on the pre-trained language model BERT, including CGExpan and ProbExpan, with ProbExpan representing the prior state-of-the-art method. Lastly, the third category involves methods based on the generative LLM, GPT-4. we devised prompt templates incorporating both positive and negative seed entities to guide the model to generate target entities.
Evaluation Metrics. The objective of ultra-fine-grained ESE is to expand entity list in descending order based on their similarity with positive seed entities. Following prior research, MAP@$K$ and P@$K$ are employed as evaluation metrics. MAP@$K$ is computed as follows:
[…]
where $Q$ stands for the set of all queries, $\mbox{AP}_{K}\left(L_{q},\mathcal{P}\right)$ denotes the average precision at position $K$ with the ranked list $L_{q}$ and ground-truth list $\mathcal{P}$. P@$K$ is the precision of the top-$K$ entities.
[…]
Table 2. Main experiment results. We report the MAP and P scores corresponding to Pos/Neg/Comb. The highest average score is bolded and the next highest average score is underlined.
Method Type Method Metric Type MAP P Avg
@10 @20 @50 @100 @10 @20 @50 @100
Probability Based SetExpan Pos $\uparrow$ 13.41 11.83 10.66 10.79 20.10 19.97 21.88 26.39 16.88
Neg $\downarrow$ 4.06 3.77 3.71 4.20 7.66 8.10 10.92 17.44 7.48
Comb $\uparrow$ 54.67 54.03 53.48 53.30 56.22 55.93 55.48 54.48 54.70
CaSE Pos $\uparrow$ 16.72 13.74 11.60 10.91 24.58 23.28 26.86 30.53 19.78
Neg $\downarrow$ 5.33 4.32 3.63 3.50 10.22 10.10 12.79 15.96 8.23
Comb $\uparrow$ 55.69 54.71 53.99 53.70 57.18 56.59 57.03 57.28 55.77
Retrieval Based CGExpan Pos $\uparrow$ 21.64 19.72 19.11 20.22 30.61 31.24 38.39 50.03 28.87
Neg $\downarrow$ 6.15 6.54 8.03 9.96 12.29 16.37 27.38 41.72 16.06
Comb $\uparrow$ 57.75 56.59 55.54 55.13 59.16 57.44 55.50 54.15 56.41
ProbExpan Pos $\uparrow$ 21.86 22.11 22.80 23.89 38.08 39.41 47.02 62.71 34.74
Neg $\downarrow$ 6.72 8.16 10.85 13.47 15.12 19.92 34.51 56.48 20.65
Comb $\uparrow$ 57.57 56.98 55.97 55.21 61.48 59.75 56.25 53.12 57.04
Generation Based GPT4 Pos $\uparrow$ 37.20 35.37 35.49 35.59 47.12 48.87 55.31 62.22 44.65
Neg $\downarrow$ 6.04 6.61 8.03 8.35 10.40 15.06 24.57 33.63 14.09
Comb $\uparrow$ 65.58 64.38 63.73 63.62 68.36 66.90 65.37 64.29 65.28
Retrieval Based (Ours) RetExpan Pos $\uparrow$ 41.73 39.53 38.55 39.91 54.58 58.03 66.76 77.23 52.04
Neg $\downarrow$ 8.77 9.04 10.65 13.29 16.44 21.04 34.78 56.54 21.32
Comb $\uparrow$ 66.48 65.25 63.95 63.31 69.07 68.50 65.99 60.34 65.36
+ Contrast Pos $\uparrow$ 47.45 44.68 43.63 44.20 59.83 62.02 69.36 77.92 56.14
Neg $\downarrow$ 8.02 8.98 10.89 13.05 14.83 21.23 35.47 55.12 20.95
Comb $\uparrow$ 69.72 67.85 66.37 65.57 72.50 70.39 66.95 61.40 67.59
+RA Pos $\uparrow$ 44.74 42.23 40.66 40.27 58.31 61.36 66.58 73.97 53.52
Neg $\downarrow$ 5.53 5.88 6.56 7.44 11.70 16.70 25.81 37.05 14.58
Comb $\uparrow$ 69.60 68.18 67.05 66.42 73.31 72.33 70.39 68.46 69.47
Generation Based (Ours) GenExpan Pos $\uparrow$ 46.79 45.00 42.89 40.80 59.77 62.15 66.26 66.57 53.78
Neg $\downarrow$ 7.25 8.28 8.72 8.01 15.21 21.31 27.33 28.58 15.59
Comb $\uparrow$ 69.77 68.36 67.08 66.40 72.28 70.42 69.47 68.99 69.10
+ CoT Pos $\uparrow$ 50.39 47.80 43.67 40.06 62.74 64.45 64.06 60.38 54.19
Neg $\downarrow$ 7.79 9.29 8.15 6.89 15.97 22.66 23.52 21.90 14.52
Comb $\uparrow$ 71.30 69.25 67.76 66.58 73.39 70.90 70.27 69.24 69.84
+ RA Pos $\uparrow$ 48.68 46.26 43.80 41.27 61.12 62.64 66.49 65.65 54.49
Neg $\downarrow$ 7.16 8.54 8.63 7.70 14.79 21.18 26.67 26.84 15.19
Comb $\uparrow$ 70.76 68.86 67.58 66.78 73.16 70.73 69.91 69.41 69.65
6.2. Main Experiments
The results of the main experiment are presented in Table 2, from which we can observe that:
(1) Regarding the average values of the comprehensive metrics CombMAP and CombP, our proposed GenExpan achieves optimal and sub-optimal performances with the incorporation of the chain-of-thought reasoning strategy and the entity-based retrieval augmentation strategy, respectively. Although the backbone model of GenExpan, LLaMA, has only 7b parameters, it surpasses the GPT-4, which comprises more than 1T parameters (over 200 times larger than the GenExpan). We attribute this to two factors. Firstly, our LLaMA undergoes further pretraining on the provided corpus focusing on entities, thereby reinforcing the model’s ultra-fine-grained semantic understanding of entities. Secondly, GenExpan ensures that the generated entities belong to the candidate entities by constraining the decoding process through the prefix tree. This mitigates the negative impact of introducing irrelevant entities on PosMAP and PosP. It can also be seen from the results that GenExpan outperforms GPT-4 mainly in the Pos class metrics. This constraint strategy is particularly valuable for ESE in user-customized scenarios. It ensures that the entities output by GenExpan are within user’s focus domain.
(2) Our proposed RetExpan model also consistently outperforms the leading retrieval-based model, ProbExpan, and the advanced generation-based model, GPT-4, as measured by the Comb metrics. RetExpan and ProbExpan are similar in their overall frameworks but have a large performance gap. The primary reason for this difference is that RetExpan utilizes the hidden state of the trained BERT for entity representation, while ProbExpan relies on the probability distribution of candidate entities at the [MASK] token. We believe that the hidden state, as a continuous vector in the feature space, captures the semantics of entities with finer granularity. In contrast, the probability distribution, as a discrete metric in the probability space, inherently offers relatively coarser granularity due to its limited capacity to store entity information. Consequently, ProbExpan underperforms in Ultra-ESE. This insight prompts future research into devising better representations to express the ultra-fine-grained semantics of entities, such as decoupling the base semantics of entities from the ultra-fine-grained attribute semantics, similar to the Mix-of-Expert (MoE) approach, where distinct features represent different perspectives of the semantics.
The ablation experiments presented in Table 3 further validate our hypothesis regarding the factors why RetExpan and GenExpan surpass the previous state-of-the-art models. Both methods exhibit remarkable degradation after removing the corresponding modules.
(3) The simple retrieval augmentation strategy yielded remarkable results for both RetExpan and GenExpan. Particularly for RetExpan, the strategy improves the comprehensive Comb metrics by 4.11 points on average. The strategy primarily optimizes the Neg class metrics, leading to a significant reduction in the intrusion of negative target entities. However, its effect on Pos class metrics exhibits instability. For instance, the PosMAP@50 and PosMAP@100 of RetExpan+RA are instead lower compared to the original RetExpan. This phenomenon can be attributed to the fact that the extra introductions corresponding to each entity are currently fixed and do not adapt to changes in the entity context. Consequently, this increases the homogeneity of the entity embedding between irrelevant entities (neither positive nor negative target entities) and positive seed entities under the same fine-grained semantic class to some extent. Our point is further supported by the observation that RetExpan’s MAP@100 for fine-grained semantic classes grows from 82.08 to 87.17 after the introduction of retrieval augmentation strategy. This suggests that a dynamic and more fine-grained knowledge retrieval strategy needs further investigation.
(4) Among all the enhancement strategies, ultra-fine-grained contrastive learning provides the most substantial improvement in the Pos class metrics, averaging 4.10 points (from 52.04 to 56.14). Conversely, contrastive learning brings smaller gains in Neg class metrics. This discrepancy is primarily attributed to the implementation of contrastive learning, where both positive and negative target entities are pulled away from normal negative entities originating from other semantic classes to ensure underlying semantics. This dilutes the pulling away of positive and negative target entities from each other, as they are both in the denominator of the contrastive loss and are assigned the same weight.
Furthermore, we found that directly increasing the weights of negative terms formed by positive and negative target entities is ineffective. This is because positive and negative target entities are determined by GPT-4 and inevitably contain errors. Therefore, this inspires us to devise more precise contrastive data mining methods in the future to amplify the penalty for hard negative terms formed by positive and negative target entities.
(5) The conventional probability-based methods SetExpan and CaSE attain quite low scores on both NegMAP and NegP metrics, which indicates that negative target entities are not excessively introduced. However, this does not imply that these methods exhibit a high degree of negative semantic awareness. The Pos class metrics of SetExpan and CaSE are similarly low, indicating their limited understanding of fine-grained semantic classes overall. Consequently, they struggle to recall entities with the given seed entities, leading to low scores in both Pos and Neg class metrics. We further evaluated the MAP@100 of CaSE at the fine-grained semantic class level, which only reaches 21.43, significantly lower than other methods such as 82.08 of RetExpan.
(6) Without further fine-tuning and modifications to the model structure, GPT-4 achieves excellent results, which proves that generative LLMs have great potential for addressing Ultra-ESE tasks. However, GPT-4 still fails to outperform the GenExpan, which is also generative and based on smaller LLaMA-7B. Analysis of cases reveals that there are two primary problems with GPT-4. On the one hand, it performs poorly on long-tail problems, such as U.S. national monuments, mobile phone brands, which contain a considerable number of low-frequency entities with limited information available on the Internet.GPT-4 only achieves single-digit PosMAP on these semantic classes. In contrast, GenExpan performs better, benefiting from the given contextual corpus and targeted retrieval augmentation strategies. On the other hand, GPT-4 is prone to haphazardly generate non-existent entities (e.g., fake mobile phone brands), which is referred to hallucination problem in recent work(Zhang et al., 2023a; Ji et al., 2023). We are currently unable to solve this issue by simple output post-processing. While the results of GPT-4 are impressive, there remains a lot of space for improvement.
Table 3. Ablation experiments for each module of RetExpan and GenExpan. Each module contributes to the overall performance.
Method MAP Avg
@10 @20 @50 @100
RetExpan 66.48 65.25 63.95 63.31 64.75
- Entity prediction 63.94 62.27 61.32 60.48 62.00
GenExpan 69.79 68.35 67.07 66.38 67.90
- Prefix constrain 57.03 56.64 56.33 56.1 56.53
- Further pretrain 68.58 66.67 65.23 64.23 66.18
6.3. Analysis of Negative Entities
We conducted analytical experiments from various perspectives regarding negative seed entities to answer the following questions:
Whether the introduction of negative seed entities is effective for representing ultra-fine-grained semantic classes? The negative seed entity-based entity re-ranking modules were removed from RetExpan and GenExpan. Thanks to the high scalability, it was also integrated into ProbExpan. Results of the three comparison experiments are presented in Table 5. We can observe that:
(1) After discarding the input of negative samples, RetExpan and GenExpan show a consistent decrease and increase in Pos class metrics and Neg class metrics, respectively. ProbExpan, on the other hand, demonstrates a rise of 0.33 points on average in the Comb class metrics after equipping the entity re-ranking module. This strongly suggests that the introduction of negative seed entities effectively mitigates the intrusion of negative target entities while raising the ranking of positive target entities, i.e., portraying the semantic classes at a finer granularity.
(2) Further analyzing the Pos and Neg metrics, we observe that the addition of the entity re-ranking module to ProbExpan conversely resulted in a notable decline in the PosP metrics. It’s essential to clarify the distinction between the MAP and P metrics: P@K is concerned solely with the number of target entities in the top-K entity list, regardless of their positions, while MAP@K is rank-aware, with higher values indicating that target entities are positioned closer to the top of the list. Consequently, this indicates that the incorporation of negative seed entities in ProbExpan makes a mixed impact, propelling some positive target entities to higher ranks and simultaneously causing others to be ranked lower.
(3) For RetExpan, negative seed entities have a similar impact on metrics with different K values, whereas for GenExpan, negative seed entities have a greater impact on metrics with a smaller K. This may be due to the fact that RetExpan is a one-time expansion (Kushilevitz et al., 2020; Mamou et al., 2018; Yu et al., 2019) model and GenExpan is an iterative expansion (Huang et al., 2020; Rong et al., 2016; Shen et al., 2017) model, so entities expanded later (corresponding to a larger K) are more likely to deviate from the original ground-truth semantic class. During these stages, re-ranking using negative seed entities also doesn’t work. Addressing the issue of semantic drift (Curran et al., 2007; McIntosh, 2010; Shi et al., 2014) in iterative expansion models like GenExpan is a longstanding challenge in ESE field.
Table 4. Comparison experiments when positive and negative attributes are the same and different.
Method Metric Type MAP Avg
@10 @20 @50 @100
$\mathbf{\mathcal{A}^{pos}=\mathcal{A}^{neg}}$
RetExpan Pos 43.14 41.74 41.93 42.89 42.43
Neg 5.54 6.44 8.71 11.56 8.06
Comb 68.80 67.65 66.61 65.67 67.18
RetExpan + Contrast Pos 49.68 48.73 48.77 49.50 49.17
Neg 5.76 6.51 8.64 11.62 8.13
Comb 71.96 71.11 70.06 68.94 70.52
RetExpan + RA Pos 46.29 44.03 43.46 43.61 44.35
Neg 5.45 5.86 6.84 7.56 6.43
Comb 70.42 69.08 68.31 68.03 68.96
$\mathbf{\mathcal{A}^{pos}\neq\mathcal{A}^{neg}}$
RetExpan Pos 41.73 39.53 38.55 39.91 39.93
Neg 8.77 9.04 10.65 13.29 10.44
Comb 66.48 65.25 63.95 63.31 64.75
RetExpan + Contrast Pos 46.44 43.69 42.51 43.14 43.95
Neg 8.22 9.25 11.17 13.38 10.51
Comb 69.11 67.22 65.67 64.88 66.72
RetExpan + RA Pos 44.74 42.23 40.66 40.27 41.98
Neg 5.53 5.88 6.56 7.44 6.35
Comb 69.60 68.17 67.05 66.42 67.81
Table 5. Ablation experiments on entity re-ranking module with negative seed entities.
MAP P
Method Metric Type @10 @20 @50 @100 @10 @20 @50 @100 Avg
Pos 21.86 22.11 22.80 23.89 38.08 39.41 47.02 62.71 34.74
Neg 6.72 8.16 10.85 13.47 15.12 19.92 34.51 56.48 20.65
ProbExpan Comb 57.57 56.98 55.97 55.21 61.48 59.75 56.25 53.12 57.04
Pos 23.63 23.65 24.13 25.20 37.36 39.41 46.79 62.71 35.36
Neg 6.82 8.22 10.89 13.47 14.79 19.92 34.48 56.48 20.63
+ Neg Rerank Comb 58.41 57.72 56.62 55.87 61.29 59.75 56.20 53.12 57.37
Pos 1.77 1.54 1.33 1.32 -0.72 0.00 -0.23 0.00 0.63
Neg 0.10 0.07 0.04 0.00 -0.33 0.00 -0.13 0.00 -0.03
$\triangle$ Comb 0.83 0.74 0.64 0.66 -0.19 0.00 -0.05 0.00 0.33
Pos 41.73 39.53 38.55 39.91 54.58 58.03 66.76 77.23 52.04
Neg 8.77 9.04 10.65 13.29 16.44 21.04 34.78 56.54 21.32
RetExpan (Ours) Comb 66.48 65.25 63.95 63.31 69.07 68.50 65.99 60.34 65.36
Pos 40.39 38.33 37.31 38.70 54.09 58.03 66.62 77.23 51.34
Neg 9.31 9.68 11.44 13.99 16.81 21.04 35.53 56.54 21.79
- Neg Rerank Comb 65.54 64.33 62.94 62.36 68.64 68.50 65.55 60.34 64.78
Pos -1.34 -1.20 -1.24 -1.21 -0.49 0.00 -0.14 0.00 -0.70
Neg 0.54 0.64 0.79 0.70 0.37 0.00 0.75 0.00 0.47
$\triangle$ Comb -0.94 -0.92 -1.01 -0.95 -0.43 0.00 -0.44 0.00 -0.59
Pos 46.79 45.00 42.89 40.80 59.77 62.15 66.26 66.57 53.78
Neg 7.25 8.28 8.72 8.01 15.21 21.31 27.33 28.58 15.59
GenExpan (Ours) Comb 69.77 68.36 67.08 66.40 72.28 70.42 69.47 68.99 69.10
Pos 46.08 44.31 42.17 40.10 59.26 62.15 66.17 66.57 53.35
Neg 7.85 8.71 9.20 8.45 15.70 21.31 27.44 28.58 15.90
-Neg Rerank Comb 69.12 67.80 66.49 65.83 71.78 70.42 69.37 68.99 68.98
Pos -0.71 -0.69 -0.72 -0.70 -0.51 0.00 -0.09 0.00 -0.43
Neg 0.60 0.43 0.48 0.44 0.49 0.00 0.11 0.00 0.32
$\triangle$ Comb -0.65 -0.56 -0.59 -0.57 -0.50 0.00 -0.10 0.00 -0.37
What are the roles played by negative seed entities? We first assessed the efficacy of RetExpan and its enhancement strategies by comparing scenarios where positive and negative attributes are identical or different. As shown in Table 4, we can find that:
(1) When positive and negative attributes are identical, there is no overlap between positive and negative target entities, which simplifies semantic classes and results in relatively high performance. In such cases, negative seed entities play a supportive role, primarily emphasizing user-concerned attributes and mitigating ambiguity. Conversely, when positive and negative attributes are different, the potential overlap between entities satisfying these attributes poses a significant challenge to semantic understanding, leading to lower overall performance. Here, negative seed entities play a crucial role in conveying unwanted semantics and are indispensable.
(2) The benefits derived from contrastive learning are more pronounced in semantic classes where positive and negative attributes are identical. This is because, in such scenarios, there is no overlap between positive and negative target entities, making the contrastive training pairs mined based on the similarity of positive and negative seed entities more reliable.
(3) When positive and negative attributes differ, the primary focus lies in evaluating the model’s capability to comprehend the semantics expressed by negative seed entities. We observe that retrieval augmentation demonstrates more effective in this situation. This phenomenon suggests that the understanding of negative seed entities demands more additional knowledge compared to positive seed classes. This notion is further verified in the analysis of chain-of-thought reasoning in Section 6.4.
What are the differences in challenges posed by semantic classes with varying numbers of positive and negative attributes? We analyze the performance of RetExpan on semantic classes characterized by varying quantities of positive and negative attributes in Table 6. To maintain control over variables, we exclusively present semantic classes for three combinations of attribute quantity: (1,1), (1,2), and (2,1). Imposing stricter constraints on the attributes results in a reduction in the number of associated target entities, rendering it more challenging to discern these entities accurately. Consequently, a decrease in the Pos (Neg) metrics is observed, correlating with an increase in the number of positive (negative) attributes. It’s important to emphasize that although the Neg metrics decrease with more negative attribute constraints, this does not imply a reduced difficulty in distinguishing negative target entities. Rather, identifying these negative target entities requires the model to possess more comprehensive attribute information about the negative seed entities.
Table 6. Comparison experiments on semantic classes with different numbers of positive and negative attributes.
$(|\mathcal{A}^{pos}|,|\mathcal{A}^{neg}|)$ Metric Type MAP Avg
@10 @20 @50 @100
(1, 1) Pos 41.32 39.18 38.36 39.58 39.61
Neg 8.61 9.14 10.88 13.43 10.52
Comb 66.36 65.02 63.74 63.07 64.55
(1, 2) Pos 39.76 37.65 41.49 43.44 40.59
Neg 0.47 0.98 2.44 5.57 2.37
Comb 69.64 68.33 69.52 68.93 69.11
(2, 1) Pos 38.84 32.85 36.39 41.33 37.35
Neg 2.39 3.29 5.38 9.81 5.22
Comb 68.22 64.78 65.51 65.76 66.07
6.4. Analysis of Enhancement Strategy
Table 7. Ablation experiments on contrastive learning.
Method Metric Type MAP Avg
@10 @20 @50 @100
RetExpan Pos 41.73 39.53 38.55 39.91 39.93
Neg 8.77 9.04 10.65 13.29 10.43
Comb 66.48 65.25 63.95 63.31 64.75
RetExpan +Contrast Pos 47.45 44.68 43.63 44.20 44.99
Neg 8.02 8.98 10.89 13.05 10.24
Comb 69.72 67.85 66.37 65.57 67.38
- Neg from ($L_{pos}$, $L_{neg}$) Pos 46.34 43.60 42.69 43.36 43.99
Neg 8.71 9.23 11.16 13.30 10.60
Comb 68.82 67.18 65.76 65.03 66.70
- Neg from ($L_{pos}$, $\overline{L_{0}}$) & ($L_{neg}$, $\overline{L_{0}}$) Pos 46.89 44.14 43.04 43.86 44.48
Neg 8.42 9.00 11.14 13.26 10.46
Comb 69.24 67.57 65.95 65.30 67.01
- Pos from ($L_{pos}$, $L_{pos}$) & ($L_{neg}$, $L_{neg}$) Pos 46.95 43.55 42.77 43.36 44.16
Neg 8.531 9.103 10.996 13.185 10.45
Comb 69.211 67.221 65.8875 65.0875 66.85
6.4.1. Analysis of Contrastive Learning
We conducted ablation experiments on each part of the training data for ultra-fine-grained contrastive learning to explore their impact on overall performance. As illustrated in Table 7, based on the full training data, the last three rows sequentially remove the hard negative samples comprising pairs of positive and negative target entities, the normal negative samples consisting of pairs of positive and negative target entities along with entities from other semantic classes, and the positive samples composed of entity pairs within the same ultra-fine-grained semantic class.
It can be seen that hard negative samples consisting of pairs of positive and negative target entities contribute the most to the overall performance, which is in line with our intuition of proposing ultra-fine-grained contrastive learning. We set both —$L_{pos}$— and —$L_{neg}$— to 10 to mitigate the introduction of excessive noise. Consequently, it’s evident that this part of training data wields more influence when K is small. Exploring how to mitigate the noise introduced by automatic GPT-4 labeling when —$L_{pos}$— and —$L_{neg}$— increase is a promising direction for future exploration.
Additionally, we observe that normal negative samples boost the overall performance. They ensure that positive and negative target samples do not deviate excessively from each other at the fine-grained semantic level, thereby preventing semantic collapse. Nevertheless, the integration of normal negative samples compromises the efficacy of ultra-fine-grained (hard) negative samples. This arises from the fact that the presence of normal negative samples indirectly diminishes the penalization applied to the distance of hard negative samples by the loss function, as they are all in the denominator of the loss function and have the same weights. Specifically, the incorporation of normal negative samples diminishes the average enhancement of hard negative samples on the Comb metric from 67.01 - 64.75 = 2.26 (line 4 minus line 1) to 67.38 - 66.70 = 0.68 (line 2 minus line 3). Moreover, positive samples, composed of entities belonging to the same ultra-fine-grained semantic class, similarly contribute positively to the overall performance.
Table 8. Comparison experiments with different retrieval contents on retrieval augmentation.
Method Metric Type MAP Avg
@10 @20 @50 @100
RetExpan + RA (Entity Introduction) Pos 44.74 42.23 40.66 40.27 41.97
Neg 5.53 5.88 6.56 7.44 6.35
Comb 69.60 68.18 67.05 66.42 67.81
RetExpan + RA (Wikidata Attributes) Pos 40.94 38.78 37.28 37.64 38.66
Neg 6.93 6.79 7.89 9.56 7.79
Comb 67.00 66.00 64.69 64.04 65.43
RetExpan + RA (GT Attributes) Pos 45.28 43.84 43.77 45.22 44.52
Neg 6.21 6.69 7.92 8.19 7.25
Comb 69.54 68.57 67.92 68.51 68.64
GenExpan + RA (Entity Introduction) Pos 48.68 46.26 43.80 41.27 45.00
Neg 7.16 8.54 8.63 7.70 8.01
Comb 70.76 68.86 67.58 66.78 68.50
GenExpan + RA (Wikidata Attributes) Pos 48.30 45.94 43.91 41.50 44.91
Neg 7.00 8.46 8.79 7.83 8.02
Comb 70.65 68.74 67.56 66.83 68.45
GenExpan + RA (GT Attributes) Pos 49.14 46.83 43.98 42.51 45.62
Neg 6.76 8.43 8.56 7.66 7.85
Comb 71.19 69.20 67.71 67.42 68.88
6.4.2. Analysis of Retrieval Augmentation
In this subsection, we delve into the impact of additional entity knowledge retrieved from various sources on RetExpan and GenExpan, including entity introductions and attribute information extracted from Wikidata, as well as ground-truth positive and negative attribute information. The experiment results presented in Table 8 demonstrate that each of the three knowledge is beneficial for both the original RetExpan and GenExpan.
Concerning the entity attribute information sourced from Wikidata, we randomly sampled an attribute subset and converted it into text with templates. The attributes on Wikidata are intricate and encompass numerous rarely-used attributes (e.g., the attributes of the entity ”Motorola” include its ”YouTube channel ID”). This may lead to the retrieved attributes are not positive or negative attributes that the current ultra-fine-grained semantic class really cares about, thereby hindering model comprehension. Therefore, despite the high quality of Wikidata’s attribute information, its gain to the model is relatively marginal.
Meanwhile, we can easily find that the direct inclusion of ground-truth positive and negative attribute information related to ultra-fine-grained semantic classes has the most substantial gain to the model. Yet, we posit that the potential of retrieval augmentation methods has not been fully realized. Firstly, the supplementary knowledge retrieved for each entity is static across different sentences and does not adapt to the entity’s context, potentially leading to homogenized entity representations. Crafting dynamic and ultra-fine-grained retrieval strategies deserves further exploration. Secondly, simply concatenating retrieved knowledge with the original sentences is not the optimal way to leverage it. Exploring more effective methods for knowledge exploitation is also essential.
6.4.3. Analysis of Chain-of-thought Reasoning
Table 9. Comparison experiments on chain-of-thought reasoning. The impact of reasoning depth and reasoning precision on performance is investigated. GT: Ground-truth. Gen: Generated. CN: Class name. Pos/Neg: Positive/Negative attribute information.
Method Metric Type MAP Avg
@10 @20 @50 @100
GenExpan Pos 46.84 44.99 42.88 40.79 43.88
Neg 7.27 8.29 8.74 8.02 8.08
Comb 69.79 68.35 67.07 66.38 67.90
GenExpan + CoT (GT CN) Pos 49.20 46.84 43.67 41.04 45.19
Neg 7.43 9.15 9.17 8.36 8.53
Comb 70.89 68.84 67.25 66.34 68.33
GenExpan + CoT (Gen CN) Pos 49.33 46.97 43.51 40.82 45.15
Neg 7.37 8.97 8.74 7.89 8.24
Comb 70.98 69.00 67.38 66.46 68.46
GenExpan + CoT (Cen CN + Gen Pos) Pos 50.39 47.80 43.67 40.06 45.48
Neg 7.79 9.29 8.15 6.89 8.03
Comb 71.30 69.25 67.76 66.58 68.72
GenExpan + CoT (Cen CN + GT Pos) Pos 50.77 48.25 44.38 40.67 46.02
Neg 7.68 9.16 8.12 6.83 7.95
Comb 71.55 69.54 68.13 66.92 69.04
GenExpan + CoT (Cen CN + Gen Pos + Gen Neg) Pos 49.78 47.24 42.68 39.01 44.68
Neg 7.60 8.97 7.98 6.75 7.82
Comb 71.09 69.14 67.35 66.13 68.43
GenExpan + CoT (Cen CN + GT Pos + GT Neg) Pos 51.02 49.19 44.86 41.40 46.62
Neg 7.30 8.70 7.59 6.53 7.53
Comb 71.86 70.25 68.63 67.43 69.54
Experiments in Table 9 examine the chain-of-thought with varying reasoning depths and assess the impact of reasoning precision on model performance. From the results, we observe that:
(1) Deeper reasoning proves to be more advantageous for Ultra-ESE, as it facilitates a more explicit understanding of the implicit ultra-fine-grained semantics of seed entities.
(2) Surprisingly, the use of manually labeled ground-truth class names is not superior to those derived through reasoning with the LLaMA-7B. Case study of the generated class names reveal that LLaMA-7B can produce class names encapsulating positive attribute information based on the given positive seed entities. For instance, when given entities about U.S. airports with the “location” attribute, GenExpan accurately infers more specific semantic classes that reflect positive attribute information, e.g., “Airports in Michigan”. This suggests that reasoning results are not necessarily inferior to manually annotated “ground-truth” labels, as chain-of-thought reasoning induces LLMs to attend to details that human ignores.
(3) Incorporating negative attribute information obtained from LLaMA-7B’s reasoning leads to a decline in model performance. Reasoning about negative attributes poses greater challenges than positive attributes due to constraints from two aspects: first, the negative seed entities are identical on negative attributes; second, the positive seed entities pose different values on negative attributes compared to the negative seed entities. Our analysis reveals that the current LLaMA-7B model struggles to deduce the relevant negative attributes by comparing positive and negative seed entities. This difficulty is particularly pronounced for long-tailed entities (such as ancient musical instruments) and semantic classes with diverse attributes, which indicates that enhancing the ultra-fine-grained perception of entities within LLMs is a promising avenue for future research.
6.5. Explore Experiments on Different Paradigm Model Interactions
Table 10. Experiments on the interaction of RetExpan and GenExpan.
Method Metric Type MAP Avg
@10 @20 @50 @100
RetExpan Pos 41.73 39.53 38.55 39.91 39.93
Neg 8.77 9.04 10.65 13.29 10.44
Comb 66.48 65.25 63.95 63.31 64.75
RetExpan + GenExpan Pos 45.93 44.37 42.96 42.35 43.90
Neg 7.03 7.51 9.04 9.48 8.27
Comb 69.45 68.43 66.96 66.43 67.82
GenExpan Pos 46.84 44.99 42.88 40.79 43.88
Neg 7.27 8.29 8.74 8.02 8.08
Comb 69.79 68.35 67.07 66.38 67.90
GenExpan + RetExpan Pos 48.91 46.09 44.13 42.49 45.41
Neg 6.24 6.80 7.56 7.84 7.11
Comb 71.33 69.64 68.29 67.32 69.15
We are curious whether retrieval-based and generation-based frameworks can interact and mutually reinforce each other. Actually, in the process of enhancing the retrieval-based RetExpan, we have leveraged the generative LLMs GPT-4. At the same time, when enhancing the generation-based GenExpan, we have utilized the retrieval-based approach to retrieve extra entity knowledge. In this section, we explore the facilitative relationship between the two paradigms at a higher level. Specifically, for the A+B model in Table 10, we first allow Model A to obtain 1000 entities from the complete candidate entities, which guarantees high entity recall. Then, model B executes on the obtained 1000 entities. To our surprise, no matter whether RetExpan or GenExpan is deployed first, both of them enhance the performance on all metrics considerably. These two different paradigm frameworks have their own advantages. The retrieval-based framework measures the semantic relation between candidate and seed entities by feature similarity, which inherently fits the nature of the retrieval task. Thus only by utilizing smaller models, retrieval-based frameworks can achieve competitive results, and be easily migrated to specific domains. Generation-based frameworks, on the other hand, typically rely on LLMs, which offer a broader knowledge base and greater semantic awareness, but the LLMs tend to be less efficient. Therefore, how to make the two types of models learn collaboratively, promote each other, and then achieve automatic evolution is a promising direction for solving Ultra-ESE.
Appendix A Prompts
When constructing contrastive learning training data, we use GPT’s classification to obtain $L_{pos}$ and $L_{neg}$. For each entity in the top-$T$ of $L_{0}$, determine whether its attributes are consistent with $S^{pos}$ ($S^{neg}$). If it is consistent, GPT should output 1; otherwise, GPT should output 0. Next, these entities that GPT deems consistent with the attributes of $S^{pos}$ ($S^{neg}$) will be merged with $S^{pos}$ ($S^{neg}$) to form $L_{pos}$ ($L_{neg}$). The prompt to classify entity is shown in Table 13.
In GenExpan, each round uses three example entities to construct $Prompt_{g}$, which is used to generate entities that are semantically similar to them. $Prompt_{g}$ is as shown in Table 14.
In the process of enhancing GenExpan with Chain of Thought prompting, we employ LLM to generate fine-grained class names. The prompt, $Prompt_{c}$, is shown in Table 15.
Appendix B Implementation Details of RetExpan
In RetExpan, sentences are tokenized using the WordPiece tokenizer and then fed into a 12-layer Transformer initialized with $BERT_{BASE}$ weights. To optimize training efficiency and preserve semantic knowledge learned by BERT, we freeze the first 11 layers of the encode, thus only the last layer is fine-tuned. When tokenizing, to ensure that mentions of entities in long sentences are not truncated, we have implemented entity focus, guaranteeing the presence of entities within the sentence. During expansion, we use one-shot expansion to directly obtain the preliminary expansion results $L_{0}$. In RetExpan with contrastive learning strategy, each epoch during training alternates between computing and optimizing entity prediction loss and contrastive learning loss.
To ensure that the encoder acquires knowledge from the corpus, we trained it for 20 epochs on 8 RTX 3090 GPUs. The hyperparameters for training learning rate, batch size, weight decay, label smoothing factor $\eta$ were set to 4e-5, 128, 1e-2, and 0.075 respectively.
Appendix C Implementation Details of GenExpan
In GenExpan, we use LLaMA-7b as the base model. Initially, we train for 1 epoch on the corpus using 6 A100 GPUs. The training hyperparameters learning rate, batch size, gradient accumulation steps, weight decay, gradient clipping are set to 1e-5, 4, 8, 1e-4, and 1.0 respectively.
During entity generation in GenExpan, we utilize prefix-constrained beam search with a beam size of 40 to generate 40 entities at a round. For entity selection, we compute the positive similarity score for each entity and select those whose scores are in the top 0.7 as the results for the current round. When no new entity is generated for 20 consecutive rounds, the generation process will end and move to re-ranking.
Appendix D Details of UltraWiki
UltraWiki contains 10 fine-grained semantic classes, which fall under 5 coarse-grained semantic categories: Organization, Location, Product, Person, and Miscellaneous. The number of entities in each fine-grained semantic class ranges from 45 to 952. Each fine-grained semantic class has 2 to 3 attributes, and depending on the combination of attributes, several ultra-fine-grained semantic classes can be derived from a single fine-grained class. Detailed data is displayed in Table 11.
The number of combinations from positive and negative attributes varies significantly, leading to substantial differences in the number of ultra-fine-grained semantic classes. The count of UltraWiki’s positive and negative attributes and their corresponding ultra-fine-grained classes are presented in Table 12.
Figure 7. Parameter analysis experiments.
Table 11. Fine-grained semantic classes detail. CLS.: Semantic Class
Coarse CLS. Fine-grained CLS. #Entities #Ultra-fine-grained CLS. Attributes
Organization Canada universities 99 10 ¡Loc-Province¿, ¡Type¿
Location China cities 675 50 ¡Province¿, ¡Prefecture¿
Countries 190 68 ¡Continent¿, ¡Driving-Side¿, ¡Per-Capita-Income¿
US airports 370 74 ¡Role¿, ¡Loc-State¿
US national monuments 112 12 ¡Loc-State¿ ¡Agency¿
Product Mobile phone brands 159 7 ¡Loc-Continent¿, ¡Status¿
Percussion instruments 128 10 ¡Type¿, ¡Source-Continent¿
Person Nobel laureates 952 11 ¡Prize¿, ¡Gender¿
US presidents 45 5 ¡Party¿, ¡Birth-State¿
Miscellaneous Chemical elements 118 14 ¡Period¿, ¡Phase-at-R.T.¿
Table 12. Types of ultra-fine-grained semantic classes. CLS.: Semantic Class
$|\mathcal{A}^{pos}|$ $|\mathcal{A}^{neg}|$Ultra-fine-grained CLS.
1 1 238
1 2 5
2 1 9
2 2 7
3 3 2
Appendix E Parameter analysis
To explore the sensitivity of each key hyperparameter in RetExpan and GenExpan, we conducted a comprehensive analysis of the smoothing factor $\eta$ and the re-ranking segment length $l$ in RetExpan, as well as the number of positive and negative entities mined in the contrastive learning strategy $|L_{pos}|$ and $|L_{neg}|$, the Top-$p$ of entity selection module in GenExpan, and the re-ranking segment length $l$, respectively. The results are depicted in Figure 7, from which we can observe that:
(1) The label smoothing $\eta$ serves to mitigate the penalty for entities possessing similar semantics to ground-truth entities. In Ultra-ESE, there are fewer entities exhibiting similarity to ground-truth entities at the ultra-fine-grained semantic level, thereby diminishing the utility of label smoothing. Consequently, the fluctuation in model performance due to parameter variations is relatively not significant.
(2) Excessive segment length $l$ in both RetExpan and GenExpan equates to degrading the segmented re-ranking strategy to a naive re-ranking approach, thereby introducing noise from irrelevant entities, as mentioned in the main text. Hence, there is a general decrease in both PosMAP@K as the segment length increases, with larger K values leading to greater declines.
(3) Regarding the ultra-fine-grained contrastive learning strategy in RetExpan, we observe that mining too many or too few positive and negative entities using GPT-4 results in performance degradation. Insufficient positive and negative entities fail to fully stimulate the potential of contrast learning, whereas an excess of such entities introduces noisy entities. Therefore, as the number of entities to be mined increases gradually, PosMAP initially rises and then falls, and NegMAP exhibits a slight initial decline followed by a rise. However, overall, these two parameters demonstrate robustness, with even the worst-performing set of values outperforming the original RetExpan.
(4) A trade-off is also necessary for the Top-$p$. A too small Top-$p$ may fail to ensure the diversity of newly generated entities, while a too large Top-$p$ may introduce an excessive number of irrelevant entities during iterative expansion. Consequently, we observe a decrease in PosMAP and a notable increase in NegMAP when Top-p is increased from 80% to 90%, which indicates that there is considerable space for improving the current rejection strategy for negatively targeted entities.
Appendix F Analysis of Model Size
We utilize various families and sizes of LLMs as backbone model of GenExpan, including 560M, 1B7, 3B, 7B1 for BLOOM, as well as 7B and 13B for LLaMA. The experiment results depicted in Figure 8 align well with the expectation. Both the BLOOM and LLaMA families satisfy the scaling law: larger models tend to be more effective in Ultra-ESE, yielding better results. Moreover, LLaMA-7B outperforms BLOOM-7B1 at the same scale.
Figure 8. Comparison experiments on different LLM families and sizes.
Figure 9. Case studies of GenExpan and the retrieval augmentation, chain-of-thought reasoning techniques. $+$ and - - - stand for positive and negative target entities, respectively. !!! represents irrelevant entities belonging to the same fine-grained semantic class as the seed entities.
Appendix G Case Study
Figure 9 presents some case studies of GenExpan and the corresponding retrieval augmentation, chain-of-thought reasoning techniques. Overall, it is evident that GenExpan predominantly avoids expanding entities that do not belong to the same semantic class as the positive and negative seed entities. This emphasizes the relative tractability of traditional ESE tasks compared to the significant challenges posed by Ultra-ESE for existing LLMs, where the model cannot exclude the expansion of negative target entities and irrelevant entities (belonging to the same fine-grained semantic class).
For retrieval augmentation, more positive target entities are introduced after fetching the additional knowledge. However, for some negative target entities that fulfill positive attributes, the retrieved knowledge may contain the introduction about its positive attributes, conversely misleading the model to overlook its negative attributes. For example, for the negative target entity “Xinxiang” in the second column, it also belongs to Henan province, i.e., it satisfies the requirement of positive attributes. Therefore, after concatenating the retrieved entity introduction “Xinxiang is a prefecture-level city in northern Henan province, China.”, it is incorrectly introduced instead.
For chain-of-thought reasoning, an interesting example is shown in Figure 9. We find that the ground-truth positive target semantics is “low-income countries”, but the model erroneously infers the positive semantic class as “African countries”, which encompasses relatively more low-income countries, and thus improves the recall of the positive target entities. Of course, some high-income African countries such as “Seychelles” are also incorrectly introduced. In fact, this error could have been avoided since the provided positive seed entity contains the non-African country “Saint Vincent and the Grenadines”. This also shows the potential for improvement in the chain-of-thought reasoning strategy.
Table 13. The prompt used to select the entities that are consistent with the attributes of $S^{pos}$ ($S^{neg}$) in the top-$T$ entities of $L_{0}$ to construct $L_{pos}$ ($L_{neg}$).
I have a task that involves classifying candidate entities based on their alignment with a seed entity set. The seed entities are grouped together because they share certain attributes, referred to as seed attributes. I will provide a list of seed entities along with their seed attributes. Additionally, I have a list of candidate entities that are similar to the seed entities but may not necessarily share the same seed attributes. I need you to identify the seed attributes and use them to classify each candidate entity into one of two categories: 1) consistent with the seed entity set in terms of seed attributes, or 0) inconsistent with the seed entity set in terms of seed attributes. For the given N candidate entities, please output N values, each being 1 or 0, indicating whether each candidate is consistent (1) or inconsistent (0) with the seed entity set based on the seed attributes.
Input:
Seed entities: [Mark Twain, Ernest Hemingway, F. Scott Fitzgerald]
Candidate entities: [J.K. Rowling, Stephen King, Agatha Christie, John Steinbeck, Harper Lee, Charles Dickens, Virginia Woolf], total 7 entities
Output:
”result”: [0,1,0,1,1,0,0]
Input:
Seed entities: [Golden Retriever, German Shepherd, Labrador Retriever]
Candidate entities: [Bengal Tiger, Beagle, Siberian Husky, African Elephant, Pug], total 5 entities
Output:
”result”: [0,1,1,0,1]
Input:
Seed entities: [{Entity1}, {Entity2}, {Entity3}]
Candidate entities: [{Entity1’}, {Entity2’}, {Entity3’}, …], total {} entities
Output:
Table 14. The prompt used to generate entities that are semantically similar to 3 given entities.
iron, copper, aluminum and zinc.
math, physics, chemistry and biology.
{Entity1}, {Entity2}, {Entity3} and ______
Table 15. The prompt used to generate a class name that covers the given entities.
Generate a class name that accurately represents the following entities. This class name should encompass all the given entities and reflect their shared characteristics.
Examples:
[[[[Tiger, Lion, Cheetah]]]] $\rightarrow$ Big Cats
[[[[Shakespeare, Tolstoy, Hemingway]]]] $\rightarrow$ Famous Authors
[[[[Mercury, Venus, Mars]]]] $\rightarrow$ Planets in the Solar System
[…]

### v3

@@
 UltraWiki: Ultra-fine-grained Entity Set Expansion with Negative Seed Entities
-Abstract.
-Entity Set Expansion (ESE) aims to identify new entities belonging to the same semantic class as a given set of seed entities.
-Traditional methods primarily relied on positive seed entities to represent a target semantic class, which poses challenge for the representation of ultra-fine-grained semantic classes.
-Ultra-fine-grained semantic classes are defined based on fine-grained semantic classes with more specific attribute constraints.
-Describing it with positive seed entities alone cause two issues: (i) Ambiguity among ultra-fine-grained semantic classes.
-(ii) Inability to define “unwanted” semantic.
-Due to these inherent shortcomings, previous methods struggle to address the ultra-fine-grained ESE (Ultra-ESE).
-To solve this issue, we first introduce negative seed entities in the inputs, which belong to the same fine-grained semantic class as the positive seed entities but differ in certain attributes.
-Negative seed entities eliminate the semantic ambiguity by contrast between positive and negative attributes.
-Meanwhile, it provide a straightforward way to express “unwanted”.
-To assess model performance in Ultra-ESE and facilitate further research, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE.
-UltraWiki encompasses 50,973 entities and 394,097 sentences, alongside 236 ultra-fine-grained semantic classes, where each query of them is represented with 3-5 positive and negative seed entities.
-A retrieval-based framework RetExpan and a generation-based framework GenExpan are proposed to comprehensively assess the efficacy of large language models from two different paradigms in Ultra-ESE.
-Moreover, we devised three strategies to enhance models’ comprehension of ultra-fine-grained entities semantics: contrastive learning, retrieval augmentation, and chain-of-thought reasoning.
+Abstract
+Entity Set Expansion (ESE) aims to identify new entities belonging to the same semantic class as the given set of seed entities.
+Traditional methods solely relied on positive seed entities to represent the target fine-grained semantic class, rendering them tough to represent ultra-fine-grained semantic classes.
+Specifically, merely relying on positive seed entities leads to two inherent shortcomings: (i) Ambiguity among ultra-fine-grained semantic classes.
+(ii) Inability to define “unwanted” semantics.
+Hence, previous ESE methods struggle to address the ultra-fine-grained ESE (Ultra-ESE) task.
+To solve this issue, we first introduce negative seed entities in the inputs, which jointly describe the ultra-fine-grained semantic class with positive seed entities.
+Negative seed entities eliminate the semantic ambiguity by providing a contrast between positive and negative attributes.
+Meanwhile, it provides a straightforward way to express “unwanted”.
+To assess model performance in Ultra-ESE and facilitate further research, we also constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE.
+UltraWiki encompasses 50,973 entities and 394,097 sentences, alongside 236 ultra-fine-grained semantic classes, where each class is represented with 3-5 positive and negative seed entities.
+Moreover, a retrieval-based framework RetExpan and a generation-based framework GenExpan are proposed to provide powerful baselines for Ultra-ESE.
+Additionally, we devised two strategies to enhance models’ comprehension of ultra-fine-grained entities’ semantics: contrastive learning and chain-of-thought reasoning.
 Extensive experiments confirm the effectiveness of our proposed strategies and also reveal that there remains a large space for improvement in Ultra-ESE.
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at https://github.com/THUKElab/UltraWiki.
-1. Introduction
-Entity Set Expansion (ESE) is a critical task aiming to identify new entities belonging to the same semantic class as a given set of seed entities (Zhang et al., 2020; Li et al., 2022a, 2023d, 2023g).
-ESE plays an important role for a wide range of user-tailored applications (Li et al., 2023b; Yu et al., 2023; Li et al., 2022c; Zhang et al., 2023b), such as recommendation systems (Jacucci et al., 2021; Ma et al., 2022; Ye et al., 2023), knowledge base population (Zhu et al., 2023; Velardi et al., 2013; Li et al., 2023a, f, e) and domain text analysis (Kohita et al., 2020; Liu et al., 2022; Li et al., 2023c, 2022b), thus demanding a high degree of semantic class granularity.
-Traditional methods of ESE primarily relied on positive seed entities to represent a target semantic class.
-However, solely leveraging positive seed entities poses a significant challenge in representing ultra-fine-grained semantic classes.
-Fine-grained semantic classes are concrete at the concept level, e.g., mobile phone brands, car brands, and the target entities of different semantic classes are typically quite different.
-In contrast, ultra-fine-grained semantic classes are defined based on concepts with more specific attribute constraints, such as mobile phone brands using Android, Asian mobile phone brands.
-Notably, ultra-fine-grained semantic classes within the same fine-grained class often demonstrate significant overlap in target entities.
-As shown in Figure 1, describing the target semantic class merely with positive seed entities cause two issues: (i) Ambiguity among ultra-fine-grained semantic classes.
-For instance, distinguishing between Mobile phone brands using Android and American Mobile phone brands using only positive seeds like {Motorola, Microsoft Mobile, Google} may lead to ambiguous expansion results.
-These seeds do not capture the distinct features that delineate each class.
+All the codes, dataset, and supplementary notes are available at https://github.com/THUKElab/UltraWiki.
+Index Terms:
+Entity Set Expansion, Knowledge Mining, Language Model, Ultra-fine-grained Semantic Understanding
+I Introduction
+Figure 1: The figure illustrates how relying only on positive seeds can lead to ambiguity between similar classes (e.g., Mobile phone brands using Android vs.
+American Mobile phone brands) and the inability to effectively define negative constraints (e.g., Mobile phone brands not made by Asian manufacturers).
+The introduction of negative seed entities in the input solves these issues.
+Entity Set Expansion (ESE) is a critical task aiming to expand new entities belonging to the same semantic class as the given set of seed entities [1, 2, 3, 4].
+For example, given the input seed entity set {Motorola, Xiaomi, Nokia}, an ESE model will output more new entities (e.g., “Huawei”, “Samsung”, …) that all belong to the same Mobile Phone brands class as the seed entities.
+ESE plays an important role for a wide range of user-tailored applications [5, 6, 7, 8, 9], thus demanding a high degree of semantic class granularity.
+For instance, in the context of recommendation systems [10, 11], a more nuanced ESE method can benefit more precise product recommendations.
+Conventionally, existing ESE methods focus on the expansion of “wanted” fine-grained semantic classes merely based on positive seed entities, such as Car Brand and the above mentioned Phone Brand.
+However, current ESE methods struggle to expand ultra-fine-grained semantic classes [12, 13, 14], which involve more specific attribute constraints (e.g., mobile phone brands using Android, Asian mobile phone brands).
+Since ultra-fine-grained semantic classes from the same concept often exhibit substantial overlap in target entities, rendering them tough to be represented only with a handful of positive seed entities.
+Specifically, as shown in Figure 1, merely relying on positive seed entities to represent ultra-fine-grained semantic classes causes two issues: (i) Ambiguity among ultra-fine-grained semantic classes.
+For instance, with the positive seed entities {Motorola, Microsoft Mobile, Google}, it is confusing whether the user intends to extend Mobile phone brands using Android or American Mobile phone brands.
+Since the positive seed entities alone are insufficient to delineate the distinct class features.
 (ii) Inability to define “unwanted”.
-Positive seeds alone cannot represent the “unwanted” semantic.
-Considering when users wish to find Mobile phone brands not made by Asian manufacturers, it is impractical to define this semantic class by enumerating phone brands from other continents (Europe, North America, Africa, etc.) in the positive seeds.
-Since the size of the positive seeds is limited, and it’s also unreasonable to expect users to specify all potential attribute values.
-Figure 1. Semantic classes at different levels of granularities, and issues with positive seed entity only.
-Previous exploratory works mainly made efforts towards fine-grained expansion.
-ProbExpan (Li et al., 2022a) refines fine-grained semantic boundaries by mining in-batch hard negative samples.
-FGExpan (Xiao et al., 2023) guides the expansion process to yield more precise results for the fine-grained semantic class with the existing taxonomy.
-MultiExpan (Li et al., 2023g) leverages the visual information of entities to differentiate between semantic classes.
-Nonetheless, these methods primarily concentrate on conceptual level distinctions (Although these concepts are somehow hard to distinguish, e.g., US states vs.
-US cities).
-They are insufficient for distinguishing ultra-fine-grained semantic classes, due to the inherent shortcoming of representing semantic classes with positive seeds only.
-Inspired by exploiting the negative feedback to improve the modeling of user purchase intent in the recommendation system (Zhao et al., 2018; Wu et al., 2020; Xie et al., 2021), we solve ultra-fine-grained ESE (Ultra-ESE) by introducing negative seed entities in the inputs.
-Specifically, the input negative seed entity belongs to the same fine-grained semantic class as the positive seed entity but differs in certain attributes.
-Positive and negative seed entities jointly represent and constrain target semantic classes, which addresses the representation granularity challenge head-on: Firstly, the contrast between positive and negative seeds on attributes can express the attribute of user interests, eliminating the semantic class ambiguity caused by positive seeds alone.
-For instance, for the aforementioned example, when the negative seeds are entities of Mobile phone brands using iOS, it indicates the user’s concern about the operation system, irrespective of the manufacturer.
-Secondly, negative seeds provide a straightforward way to express “unwanted” semantics.
-Regrettably, current ESE dataset (Shen et al., 2017; Pradhan et al., 2013) lacks the concept of negative seed entities, let alone ultra-fine-grained semantic classes.
+Positive seed entities alone fail to represent the “unwanted” semantic.
+Considering when users wish to find more Mobile phone brands not made by Asian manufacturers entities, it is impractical to define this semantic class by enumerating brands from other continents in positive seed entities.
+As the size of the positive seed entities is limited, it’s also unreasonable to expect users to specify all potentially desired attribute values.
+Due to the inherent shortcomings of the task inputs, prior ESE methods failed to conduct ultra-fine-grained expansion.
+To address this issue, inspired by exploiting negative feedback to model user intent in recommendation system [15, 16, 17], we propose enhancing ESE by incorporating negative seed entities, aiming at ultra-fine-grained ESE (Ultra-ESE).
+The input negative seed entities belong to the same fine-grained semantic class as the positive seed entities but differs in certain attributes.
+They jointly represent target semantic class, addressing the semantic granularity issue head-on: First, the contrast between positive and negative seed entities in terms of attributes highlights the user’s specific interests and eliminates the ambiguity arising from positive seed entities alone.
+For the aforementioned example, if negative seed entities are about Mobile phone brands not using Android, it indicates that the user is focused on operating system rather than manufacturer.
+Second, negative seed entities constrain the expansion space and can readily express “unwanted” semantics.
+Regrettably, previous ESE dataset [18, 19] lacked the concept of negative seed entities, let alone ultra-fine-grained semantic classes.
 To bridge this gap, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE.
-Derived from Wikipedia, UltraWiki encompasses 10 fine-grained semantic classes, alongside 50,973 entities and 394,097 sentences.
-We determined 2-3 attributes for each fine-grained semantic class and annotated each entity with attribute values utilizing a semi-automatic approach.
-Based on the identified attributes, we further constructed 236 ultra-fine-grained semantic classes, in which each semantic class is represented with 3-5 positive and negative seed entities, respectively.
-In other words, there is an average of 23 (236/10-1) hard negative semantic classes for each ultra-fine-grained semantic class, which poses a significant challenge to the model’s capability for ultra-fine-grained semantic comprehension.
-Even advancing GPT-4 can’t handle it well.
-In experiments, we evaluate the performance of existing methods on UltraWiki.
-Furthermore, we designed a retrieval-based framework with Encoder-only models (BERT (Devlin et al., 2018)) and a generation-based framework with Decoder-only models (LLaMA (Touvron et al., 2023)) to comprehensively assess the efficacy of large language models (LLMs) from two different paradigms on the Ultra-ESE task.
+Derived from Wikipedia, UltraWiki encompasses 10 fine-grained semantic classes, 50,973 entities, and 394,097 sentences.
+During the curation, we annotated 2-3 attributes for each fine-grained semantic class to further construct ultra-fine-grained semantic classes.
+Based on the permutations of attributes, 235 ultra-fine-grained semantic classes are constructed, with each semantic class represented with 3-5 positive and negative seed entities.
+On average, each ultra-fine-grained semantic class contains 23 (236/10-1) hard negative semantic classes, which belong to the same fine-grained semantic class and may exhibit substantial overlap in target entities.
+Experiments proved that even advancing GPT-4 can not handle it well.
+In experiments, we evaluated existing ESE methods on UltraWiki.
+Furthermore, to comprehensively assess the efficacy of large language models (LLMs) with two different paradigms on the Ultra-ESE task, we designed a retrieval-based framework, RetExpan, with encoder-only LLM BERT [20] and a generation-based framework, GenExpan, with decoder-only LLM LLaMA [21].
 Meanwhile, we proposed three strategies to enhance models’ ability to comprehend ultra-fine-grained semantics of entities: contrastive learning, retrieval augmentation, and chain-of-thought reasoning.
-In summary, the main contributions include:
-  • We proposed the more challenging ultra-fine-grained ESE task, and introduce negative seed entities to represent ultra-fine-grained semantic classes more precisely for the first tiem.
+In summary, the main contributions are listed as follows:
+  • We proposed the more challenging ultra-fine-grained ESE task, and incorporated negative seed entities to represent ultra-fine-grained semantic classes more precisely for the first time.
   • We constructed the first large-scale dataset UltraWiki, tailored for ultra-fine-grained ESE tasks. It encompasses 10 fine-grained semantic classes and 261 ultra-fine-grained semantic classes.
   • We designed both retrieval-based and generation-based frameworks to assess the efficacy of BERT-like and GPT-like LLMs on the UltraWiki dataset.
-Furthermore, three strategies are proposed for refining the semantic comprehension of ultra-fine-grained entities.
-  • Extensive experiments confirm the effectiveness of our proposed strategies and also reveal significant potential for enhancing ultra-fine-grained semantic comprehension of entities by LLMs.
-2. Related Work
-2.1. Methods of ESE
+Furthermore, two strategies were proposed for refining the semantic comprehension of ultra-fine-grained entities.
+  • Extensive experiments confirmed the effectiveness of our proposed strategies and also reveal significant potential for enhancing ultra-fine-grained semantic comprehension of entities by LLMs.
+II Related Work
+II-A Methods of ESE
 Recently, increasing emphasis has been placed on the expansion of fine-grained (e.g., US city) rather than coarse-grained (e.g., Location) semantic classes.
-ProbExpan (Li et al., 2022a) employs heuristic methods to mine hard negative entities of target semantic classes, thereby refining the semantic boundaries of fine-grained classes through contrastive learning (Chen et al., 2020; Gao et al., 2021).
-FGExpan (Xiao et al., 2023) leverages the existing taxonomy to direct BERT in reasoning about more fine-grained names of semantic classes.
-Besides, MultiExpan (Li et al., 2023g) integrates multimodal pre-trained models to encode visual information (images) associated with entities, which benefits the differentiation of various fine-grained semantic classes.
+ProbExpan [3] employs heuristic methods to mine hard negative entities of target semantic classes, thereby refining the semantic boundaries of fine-grained classes through contrastive learning [22, 23].
+FGExpan [24] leverages the existing taxonomy to direct BERT in reasoning about more fine-grained names of semantic classes.
+Besides, MultiExpan [4] integrates multimodal pre-trained models to encode visual information (images) associated with entities, which benefits the differentiation of various fine-grained semantic classes.
 Nevertheless, these methods do not essentially tackle the issue that solely using positive seed entities fails to sufficiently represent ultra-fine-grained semantic classes.
 In contrast, our research introduces the novel concept of incorporating negative seed entities alongside positive ones to describe an ultra-fine-grained semantic class.
-It is worth mentioning that there has been some work to incorporate negative entities (Curran et al., 2007; Jindal and Roth, 2011; Shi et al., 2014; Li et al., 2022a), though the usage of negative entities in these methods is relatively naive.
+It is worth mentioning that there has been some work to incorporate negative entities [25, 26, 27, 3], though the usage of negative entities in these methods is relatively naive.
 We point out that the role of the negative entities used in previous work is fundamentally different from ours.
-Negative entities in previous work are purely used to help determine the boundary of the target set described by positive seeds.
-In contrast, negative entities in our model are used to describe the target ultra-fine-grained classes that cannot be characterized by positive seeds alone.
-2.2. Data Resources of ESE
+Negative entities in previous work are purely used to help determine the boundary of the target set described by positive seed entities.
+In contrast, negative entities in our model are used to describe the target ultra-fine-grained classes that cannot be characterized by positive seed entities alone.
+II-B Data Resources of ESE
 Queries in the existing ESE dataset merely consist of positive seed entities.
@@
 Consequently, existing datasets lack ultra-fine-grained semantic classes.
-For instance, Wiki (Shen et al., 2017) and APR (Shen et al., 2017), derived from the Wikipedia and Reuters corpora respectively, encompass merely 3 and 8 fine-grained semantic classes.
+For instance, Wiki [18] and APR [18], derived from the Wikipedia and Reuters corpora respectively, encompass merely 3 and 8 fine-grained semantic classes.
 Some named entity recognition datasets are directly used as ESE evaluation benchmarks, they contain semantic classes with coarser granularity.
-For example, the OntoNotes (Pradhan et al., 2013) and CoNLL (Sang and De Meulder, 2003) datasets include semantic classes like organization and location.
-In contrast, UltraWiki encompasses 261 ultra-fine-grained semantic classes, which surpasses existing ESE datasets in term of both granularity and quantity of semantic classes.
-Figure 2. llustration of the UltraWiki dataset construction process.
-Figure 3. Distribution of positive and negative target entities in two different cases.
-3. Task Formulation
-Our study focuses on the ultra-fine-grained ESE task, whose input incorporates three components: query $S$, candidate entity vocabulary $V$, and corpus $D$.
-The query $S=\{S^{pos}\cup S^{neg}\}$ is a composite set comprising positive seed entities $S^{pos}=\{e^{pos}_{1},\ldots,e^{pos}_{k}\}$ and negative seed entities $S^{neg}=\{e^{neg}_{1},\ldots,e^{neg}_{k}\}$, which jointly represent a target ultra-fine-grained semantic class.
-Both $S^{pos}$ and $S^{neg}$ belong to the same fine-grained semantic class.
-The entities in $S^{pos}$ and $S^{neg}$ share same values for certain attributes $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$, respectively.
-Notably, as shown in Figure3, $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$ are not limited in quantity and may overlap or differ.
-The vocabulary $V$ enumerates the candidate entities.
-Additionally, the corpus $D$ supplies contextual information $\{e_{i},s^{i}_{1},\ldots,s^{i}_{n}\}$ for each entity $e_{i}$, in which $s^{i}_{n}$ is a sentence containing entity $e_{i}$.
-The ultimate objective of ultra-fine-grained ESE is to identify positive target entities from the vocabulary $V$ that belong to the same fine-grained semantic class, and share the same attribute values with $S^{pos}$ while distinct from $S^{neg}$.
-In an ideal entity feature space, entities that share more same attribute values shall be positioned closer.
-4. UltraWiki Dataset
-Considering the lack of negative seed entities and ultra-fine-grained semantic classes in existing ESE datasets, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE.
+For example, the OntoNotes [19] and CoNLL [28] datasets include semantic classes like organization and location.
+In contrast, UltraWiki encompasses 261 ultra-fine-grained semantic classes, which surpasses existing ESE datasets in terms of both granularity and quantity of semantic classes.
+III Task Formulation
+We focus on the ultra-fine-grained ESE task, whose input consists of three components: query $S$, candidate entities $V$, and corpus $D$.
+The query $S=\{S^{pos}\cup S^{neg}\}$ is a composite set comprising positive seed entities $S^{pos}=\{e^{pos}_{1},\ldots,e^{pos}_{k}\}$ and negative seed entities $S^{neg}=\{e^{neg}_{1},\ldots,e^{neg}_{k}\}$, which belong to the same fine-grained semantic class $c$.
+The entities in $S^{pos}$ share the same values for positive attribute set $\mathcal{A}^{pos}$:
+$e.\operatorname{get}(a_{i})\equiv k_{i},\forall e\in S^{pos},\forall a_{i}\in\mathcal{A}^{pos}=\left\{\left(a_{i},k_{i}\right)\right\}_{i=1}^{|\mathcal{A}^{pos}|}$ (1)
+where $a_{i}$ and $k_{i}$ are attribute and value, respectively.
+We define positive target entities $\mathcal{P}$ as entities that share the same attribute values with positive seed entities in $\mathcal{A}^{pos}$.
+Likewise, negative target entities in $\mathcal{N}$ share the same attribute values with negative seed entities in $\mathcal{A}^{neg}$.
+Overall, the target ultra-fine-grained semantic class means class $c$ that holds the same values with positive seed entities in $\mathcal{A}^{pos}$, but different values with negative seed entities in $\mathcal{A}^{neg}$.
+Examples can be found in Figure 2.
+Hence, the ultimate objective of ultra-fine-grained ESE is to identify entities from candidate entities $V$ that belong to the $\mathcal{P}$, while being distinct from $\mathcal{N}$ (e.g., $\mathcal{P}-\mathcal{N}$).
+In an ideal feature space, entities that share more of the same attribute values should be positioned closer.
+Figure 2: Left target semantic class: mobile phone brand with Android OS, not produced by Asian manufacturer.
+Right target semantic class: mobile phone brand with Android OS.
+The target entity set is $\mathcal{P}-\mathcal{N}$.
+IV UltraWiki Dataset
+Figure 3: Illustration of the UltraWiki dataset construction process. The intermediate data obtained at each step is bolded in the figure.
+Considering the lack of ultra-fine-grained semantic classes and negative seed entities in existing ESE datasets, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE.
 This section outlines the construction process of UltraWiki.
 Semantic granularity, annotation consistency, and candidate entities’ difficulty are fully considered to guarantee the quality of UltraWiki.
-4.1. Dataset Construction
-Two methods exist for constructing UltraWiki.
-A more direct method is to first collect a large number of entities and corresponding context sentences.
+IV-A Dataset Construction
+There are two strategies for constructing UltraWiki.
+A more direct strategy is to first collect a large number of entities and corresponding context sentences.
 Then, for each fine-grained semantic class, the human annotator traverses the entire entity vocabulary to identify the corresponding entity and annotates the attribute values, thereby constructing the ultra-fine-grained semantic class.
-However, despite the availability of numerous entity-centric datasets, this method is labor-intensive and prone to omissions of entities.
+However, despite the availability of numerous entity-centric datasets, this method is labor-intensive and prone to omit entities.
 Consequently, we chose a more practical strategy.
@@
 Wikipedia maintains an extensive catalog of entities associated with fine-grained semantic classes.
-We carefully chose 10 semantic classes with certain principles (discussed in Section 3.2) and then crawled the corresponding entities.
+We identified 10 fine-grained semantic classes and then crawled the corresponding entities.
 Additionally, a substantial number of entities sampled from Wikipedia pages were incorporated into the candidate entity vocabulary as negative entities.
@@
 A significant volume of text is also crawled from Wikipedia pages, where entities are uniquely identified by hyperlinks.
-Since the entities obtained in Step 1 also contain hyperlinks, we can readily establish associations between entities and the sentences containing them, thereby providing abundant contextual information about the entities.
+Since the entities obtained in Step 1 also contain hyperlinks, we can readily align entities with sentences containing them, thereby providing abundant contextual information for entities.
 Step 3.
@@
 For each fine-grained semantic class $C$, we manually select $k$ independent attributes $\mathcal{A}=\{a_{1},a_{2},\cdots,a_{k}\}$, which will be used to generate ultra-fine-grained semantic class in next step.
-Attributes of each class are shown in appendix.
+Attributes of each class are shown in supplementary notes.
 Subjective attributes are avoided like good or bad.
-We initially employed a script to automatically query Wikidata for the corresponding attribute value.
-For the remaining attributes that could not be automatically annotated, we employed professional annotators to manually label them, ensuring each entity was reviewed three times for accuracy and consistency.
+We first query Wikidata API for the corresponding attribute value with Python script.
+For the remaining attributes that could not be automatically annotated, we employed human annotators to manually label them, ensuring each entity was reviewed three times for accuracy and consistency.
 Step 4.
 Negative-aware Semantic Classes Generation.
-We devised an algorithm for automated generation of ultra-fine-grained semantic classes based on positive and negative attributes.
-Specifically, for attributes $\mathcal{A}$ of each fine-grained semantic class $C$, we sample $n^{pos}$ and $n^{neg}$ attributes as positive and negative attributes (denoted as $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$).
+We devised an algorithm for automated generation of ultra-fine-grained semantic classes based on annotated attributes.
+Specifically, for attributes $\mathcal{A}$ of each fine-grained semantic class $C$, we sample $m$ and $n$ attributes as positive and negative attribute set (i.e., $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$).
 Subsequently, we pick a value for all the attributes in $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$.
-This process yields a list of positive values $\mathcal{V}^{pos}$ and a list of negative values $\mathcal{V}^{neg}$, which jointly constrain the attributes of the target semantic class to achieve ultra-fine-grained control of the semantic class.
+This process yields a tuple of positive values $\mathcal{V}^{pos}$ and a tuple of negative values $\mathcal{V}^{neg}$, which jointly constrain the attributes of the target semantic class to achieve ultra-fine-grained control of the semantic class.
 When $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$ are the same, the role of $\mathcal{A}^{neg}$ is to emphasize attributes concerned by user and mitigate ambiguity.
 Conversely, when they differ, $\mathcal{A}^{neg}$ serves to convey unwanted semantics.
-We denote the set of entities satisfying $\mathcal{V}^{pos}$ and $\mathcal{V}^{neg}$ as positive target entity $\mathcal{P}$ and negative target entity $\mathcal{N}$ respectively, ensuring that both $|\mathcal{P}|$ and $|\mathcal{N}|$ exceed the minimum entity requirement $n_{thred}=6$.
-Figure 4. Heat map visualization of semantic class similarity.
-4.2. Dataset Analysis
+As described in the Section III, positive and negative target entity sets are denoted as $\mathcal{P}$ and $\mathcal{N}$.
+Both $|\mathcal{P}|$ and $|\mathcal{N}|$ are ensured to exceed the minimum entity requirement $n_{thred}=6$.
+Figure 4: Semantic similarity heatmap of ultra-fine-grained semantic classes.
+Rows/columns: averaged embeddings of ground-truth positive entities per class; cell colors: pairwise cosine similarities.
+Classes were proportionally sampled to 80 classes based on total size for clarity.
+IV-B Dataset Analysis
 Statistics of UltraWiki.
-UltraWiki is the first large-scale dataset to feature ultra-fine-grained semantic classes.
-It comprises a collection of 50,973 entities and 394,097 sentences sourced from Wikipedia.
-UltraWiki encompasses 10 fine-grained semantic classes that comprehensively cover five major entity types, including Organization, Location, Product, Person, and Miscellaneous.
+UltraWiki is the first large-scale dataset featuring ultra-fine-grained semantic classes.
+It comprises 50,973 entities and 394,097 sentences sourced from Wikipedia.
+10 fine-grained semantic classes are determined, that comprehensively cover five major entity types, including Organization, Location, Product, Person, and Miscellaneous.
 Leveraging the combination of positive and negative attributes, we further derived 261 ultra-fine-grained semantic classes.
 On average, each ultra-fine-grained semantic class comprises 63 positive target entities ($\mathcal{P}$) and 60 negative target entities ($\mathcal{N}$) that are not expected to be expanded.
-Each ultra-fine-grained semantic class encompasses 3 queries, with 3 to 5 positive seed entities and 3 to 5 negative seed entities.
+Each ultra-fine-grained semantic class encompasses 3 queries, with 3 to 5 positive and negative seed entities, respectively.
 The majority of the semantic classes are constrained by one positive and one negative attribute.
-Table 1. Comparison of ESE datasets.
+TABLE I: Comparison of ESE datasets.
 Wiki APR CoNLL ONs UltraWiki
@@
 The fine-grained semantic classes, entity corpus, and partial attributes of UltraWiki are automatically crawled from Wikipedia and Wikidata.
-Both sources are high-quality knowledge bases, curated by numerous domain experts, thereby significantly guaranteeing the quality of UltraWiki.
+Both sources are high-quality knowledge bases, curated by numerous domain experts, thereby firmly guaranteeing the quality of UltraWiki.
 For attributes that require manual annotation, we ensure that each value is annotated by three annotators.
-Eventually, the inter-annotator agreement measured by Fleiss’s Kappa (Fleiss, 1971) reaches 0.90, indicating satisfying consistency and accuracy of the annotation.
+Eventually, the inter-annotator agreement measured by Fleiss’s Kappa [29] reaches 0.90, indicating satisfying consistency and accuracy of the annotation.
 Difficulty of UltraWiki.
@@
 The challenge arises from the substantial entity overlap among ultra-fine-grained semantic classes (approximately 99% of ultra-fine-grained semantic classes intersect), which requires the model to comprehend the ultra-fine-grained distinctions of entities across varied contexts.
-Even advanced models like GPT-4 struggle to address this challenge (refer to the results in the Experimental Section).
-Additionally, our selected semantic classes includes a subset of long-tail entities, such as lesser-known Chinese cities, which typically lack extensive contextual knowledge.
+Even advanced models like GPT-4 struggle to address this challenge (refer to the results in the section of experiments).
+Additionally, our selected semantic classes include a subset of long-tail entities, such as lesser-known Chinese cities, which typically lack extensive contextual knowledge.
 Simultaneously, employing the BM25-based search, we incorporated entities highly similar to the target entities as hard negative entities in the candidate entity vocabulary.
 The inclusion of these entity types contributes to the difficulty of UltraWiki.
-The visualization of entity similarity in Figure 4 illustrates that each semantic class of UltraWiki possesses extremely high intra-class similarity.
-5. Methods
+The visualization of semantic similarity in Figure 4 demonstrates that each semantic class in UltraWiki exhibits remarkably high intra-class similarity, where each row/column represents the average embedding of all ground-truth positive entities within a specific ultra-fine-grained semantic class.
+Figure 5: Upper: overall frameworks of our RetExpan. Below: overall frameworks of our GenExpan.
+V Methods
 This section presents our proposed retrieval-based and generation-based frameworks to leverage both positive and negative seed entities, namely RetExpan and GenExpan.
@@
 Most prior ESE models are built upon retrieval-based frameworks, e.g., CGExpan and ProbExpan.
-The retrieval-based framework first encodes entities as feature vectors in a low-dimensional semantic space, and then measures the probability that a candidate entity belongs to the target semantic class via the entity representations similarity, which is naturally more suitable for retrieval tasks like ESE.
-Considering the recent emergence of generative LLMs, we also propose a generation-based framework to explore the potential of generative LLMs.
-Moreover, for each framework, two separate enhancement strategies are crafted to enhance the models’ capacity for ultra-fine-grained semantic comprehension of entities.
+The retrieval-based framework first encodes entities as features in low-dimensional semantic space, and then measures the probability that a candidate entity belongs to the target semantic class via entity feature similarity, which is naturally more suitable for retrieval tasks like ESE.
+With the advent of generative LLMs, we note their significant semantic comprehension and reasoning skills, which are developed through pre-training on large-scale corpora.
+These models can eliminate the intermediate step of entity embedding, allowing for a more streamlined, end-to-end Entity Set Expansion ESE process.
+Therefore, we also propose a generation-based framework to explore the potential of generative LLMs.
+Moreover, for both frameworks, enhancement strategies are crafted to enhance the models’ capacity for ultra-fine-grained semantic comprehension with negative seed entities.
 The proposed RetExpan and GenExpan provide comprehensive and strong baselines for Ultra-ESE.
-Figure 5. Overall Frameworks of RetExpan and GenExpan.
-5.1. Retrieval-based Framework: RetExpan
-5.1.1. Overall Framework
+V-A Retrieval-based Framework: RetExpan
+V-A1 Overall Framework
 RetExpan is structured into three steps: entity representation, entity expansion, and entity re-ranking.
-In the first stage, we design the entity encoder that extracts contextual feature vectors of entities in sentences.
+In the first stage, we design the entity encoder that extracts contextual features of entities in sentences.
 An entity is represented as the mean of the feature vectors derived from all sentences containing it.
-Entity prediction task is introduced to refine the entity representation.
-During the second stage, RetExpan obtains a preliminary list of target entities based on their similarity to the positive seed entities.
-Notably, negative seed entities are omitted in this process to ensure the recall of all entities satisfying fine-grained semantic classes.
-In the third stage, negative seed entities are also employed to re-rank the entity list obtained in the previous phase, reducing the ranking of entities aligning with the attributes of the negative seed entities.
+The entity prediction task is introduced to refine the entity representation.
+During the second stage, a preliminary list of target entities is acquired based on their similarity to the positive seed entities.
+Notably, negative seed entities are excluded in this process to ensure the recall of all entities satisfying fine-grained semantic classes.
+In the third stage, negative seed entities are also employed to re-rank the entity list obtained in the previous phase, reducing the ranking of entities aligning with negative attribute values.
 Entity Representation.
@@
 Initially, we replace entity mentions within a sentence with [MASK] tokens to construct the input for the encoder.
-Given a contextual sentence $T$ with masked entity mentions, we utilize the $\text{BERT}_{\text{BASE}}$, which comprises a 12-layer Transformer (Vaswani et al., 2017), to extract contextual embeddings:
-(1) $H=\{{h}_{1},{h}_{2},...,{h}_{L}\}=\text{BERT}_{\text{BASE}}(T)$
+Given a contextual sentence $T$ with masked entity mentions, we utilize the $\text{BERT}_{\text{BASE}}$, which comprises a 12-layer Transformer [30], to extract contextual embeddings:
+$H=\{{h}_{1},{h}_{2},...,{h}_{L}\}=\text{BERT}_{\text{BASE}}(T)$ (2)
 where $L$ is the length of the tokenized sentences.
 Ultimately, the contextual feature $\mathbf{h}$ of an entity is computed as the average of the contextual embedding mask $h_{[\text{MASK}]}$ at the mask position across all sentences containing it.
-Inspired by ProbExpan, entity prediction task is introduced to refine the entity representation.
+Inspired by ProbExpan, the entity prediction task is introduced to refine the entity representation.
 Concretely, a classification header $\mathbf{f}$ is appended to the entity encoder.
 After obtaining the hidden state $h_{[\text{MASK}]}$ at [MASK] position, it is transformed into the probability distribution of the masked entity among the candidate entities via MLP and SoftMax:
-(2) $\small\hat{y}=\mathbf{f}\left(h_{[\text{MASK}]}\right)=\operatorname{Softmax}\left(\operatorname{MLP}\left(h_{[\text{MASK}]}\right)\right),\hat{y}\in\mathbb{R}^{V_{e}}$
-where $V_{e}$ is the size of candidate entities. Then, cross-entropy loss with label smoothing (Szegedy et al., 2016) is applied to learn the latent semantics of entities:
-(3) $\small\begin{split}\mathcal{L}_{mask}=-\frac{1}{N}\sum_{i}^{N}\sum_{j}^{V_{e}}y_{i}[j]\cdot(1-\eta)\cdot log(\hat{y}_{i}[j])\\ +(1-y_{i}[j])\cdot\eta\cdot log(1-\hat{y}_{i}[j])\end{split}$
-in which ground-truth $y$ is a one-hot vector and $N$ is the batch size. Smoothing factor $\eta$ mitigates over-penalization for entities that exhibit similar semantic to the ground-truth entity.
+$\hat{y}=\mathbf{f}\left(h_{[\text{MASK}]}\right)=\operatorname{Softmax}\left(\operatorname{MLP}\left(h_{[\text{MASK}]}\right)\right),\hat{y}\in\mathbb{R}^{V_{e}}$ (3)
+where $V_{e}$ is the size of candidate entities. Then, cross-entropy loss with label smoothing [31] is applied to learn the latent semantics of entities:
+$\begin{split}\mathcal{L}_{mask}=-\frac{1}{N}\sum_{i}^{N}\sum_{j}^{V_{e}}y_{i}[j]\cdot(1-\eta)\cdot log(\hat{y}_{i}[j])\\ +(1-y_{i}[j])\cdot\eta\cdot log(1-\hat{y}_{i}[j])\end{split}$ (4)
+in which ground-truth $y$ is a one-hot vector and $N$ is the batch size. Smoothing factor $\eta$ mitigates over-penalization for entities that exhibit similar semantics to the ground-truth entity.
 Entity Expansion.
@@
 For each candidate entity $e$, we define its positive similarity score $sco^{pos}$ as follows:
-(4) $\small sco^{pos}(e)=\frac{1}{|{S}^{pos}|}\sum_{e^{\prime}\in{S}^{pos}}\mathrm{cos}\_\mathrm{sim}\left(h(e),h(e^{\prime})\right)$
+$sco^{pos}(e)=\frac{1}{|{S}^{pos}|}\sum_{e^{\prime}\in{S}^{pos}}\mathrm{cos}\_\mathrm{sim}\left(h(e),h(e^{\prime})\right)$ (5)
 We keep top-$K$ entities with the highest $sco^{pos}$ as the initial expansion list $L_{0}$.
@@
 This step further considers negative seed entities, excluding entities that share semantics with them (i.e., possess the same value on negative attribute).
-Analogous to Equation 4, we define the negative similarity score ${sco}^{neg}$ between candidate entity and negative seed entities.
+Analogous to Equation 5, we define the negative similarity score ${sco}^{neg}$ between the candidate entity and negative seed entities.
 However, directly re-ranking $L_{0}$ in ascending order utilizing $sco^{neg}$ introduces a significant number of noisy entities that do not belong to the same fine-grained semantic class as the seed entities.
@@
 This strategy facilitates a local fine ranking of $L_{0}$ and prevents the assignment of over-high rankings to noisy entities with quite low $sco^{neg}$ values.
-5.1.2. Strategy 1: Ultra-fine-grained Contrastive Learning
-To enhance the ultra-fine-grained semantic comprehension of entities, we devised attribute-aware contrastive learning.
-Conventional contrastive learning facilitates clearer semantic boundaries for fine-grained semantic classes by drawing the representation of the same semantic class entities closer and the representation of different semantic class entities further apart.
+V-A2 Enhancement Strategy 1: Ultra-fine-grained Contrastive Learning
+To further enhance the ultra-fine-grained semantic comprehension capability of RetExpan, we devised attribute-aware contrastive learning.
+Conventional contrastive learning [32] facilitates clearer semantic boundaries for fine-grained semantic classes by drawing the representation of the same semantic class entities closer and the representation of different semantic class entities further apart.
 Although proven to be effective, basic contrastive learning is not precise enough for Ultra-ESE as it neglects entity differences inside the same semantic class.
-For instance, ”Samsung” and ”Motorola” may serve as positive pairs in terms of the attribute ”Operating System” but as negative pairs for the attribute ”Manufacturer”.
-Ideally, entities with more attributes in common should exhibit closer distance in the entity space.
+For instance, “Samsung” and “Motorola” may serve as positive pairs in terms of the attribute “Operating System” but as negative pairs for the attribute “Manufacturer”.
+Ideally, entities with more attributes in common should exhibit closer distance in feature space.
 To this end, we construct ultra-fine-grained training data based on positive and negative seed entities to implicitly incorporate attribute-based differences into entity representations.
 Ultra-fine-grained Training Data.
-We prompt GPT-4111All specific prompts used for querying GPT-4 can be found in the Appendix.
-We use GPT-4-1106-preview.
+We prompt GPT-4111All specific prompts used for querying GPT-4 can be found in supplementary notes.
+We use GPT-4-1106.
 to deduce potential positive (negative) attributes and return T entities from the initial list $L_{0}$ that are most similar to the given positive (negative) seed entities to form the entity list $L_{pos}$ ($L_{neg}$).
@@
 Specifically, we construct the contrastive training pairs as follows:
-(5) $\small\begin{split}\mathbb{P}_{pos}=\{(x,x^{\prime})|e(x)\in L_{pos},e(x^{\prime})\in L_{pos}\}\cup\\ \{(x,x^{\prime})|e(x)\in L_{neg},e(x^{\prime})\in L_{neg}\}\cup\{(x,x^{\prime})|e(x)=e(x^{\prime})\},\end{split}$
-(6) $\small\begin{split}\mathbb{P}_{neg}=\{(x,y)|e(x)\in L_{pos},e(y)\in L_{neg}\}\cup\\ \{(x,z)|e(x)\in L_{pos}\cup L_{neg},e(z)\in\overline{L_{0}}\},\end{split}$
+$\small\begin{split}\mathbb{P}_{pos}=\{(x,x^{\prime})|e(x)\in L_{pos},e(x^{\prime})\in L_{pos}\}\cup\\ \{(x,x^{\prime})|e(x)\in L_{neg},e(x^{\prime})\in L_{neg}\}\cup\{(x,x^{\prime})|e(x)=e(x^{\prime})\},\end{split}$ (6)
+$\small\begin{split}\mathbb{P}_{neg}=\{(x,y)|e(x)\in L_{pos},e(y)\in L_{neg}\}\cup\\ \{(x,z)|e(x)\in L_{pos}\cup L_{neg},e(z)\in\overline{L_{0}}\},\end{split}$ (7)
 where $e(x)$ indicates the entity of the training sample $x$.
@@
 Contrastive Loss.
-We choose the classical InfoNCE loss (Oord et al., 2018) as the contrastive loss.
+We choose the classical InfoNCE loss [33] as the contrastive loss.
 Contrastive learning is conducted in a new hypersphere space to prevent semantic collapse, which is transformed by another MLP-based mapping head $\mathbf{f_{cl}}$ and $l$-2 normalization.
-5.1.3. Strategy 2: Entity-based Retrieval Augmentation
-The contextual information of entities is sourced from the provided corpus ${D}$, which might not offer adequate information for ultra-fine-grained semantic distinctions, particularly for long-tailed entities.
-Hence, we gather the corresponding introduction text of entities from the external corpus, Wikidata, to augment the entity knowledge.
-During both training and inference, these entity introductions serve as prefixes for all sentences, thereby enriching the model’s ultra-fine-grained comprehension of the entities.
-5.2. Generation-based Framework: GenExpan
-5.2.1. Overall Framework
+V-B Generation-based Framework: GenExpan
+V-B1 Overall Framework
 To explore the potential of emerging generative LLMs, we devised the generative framework GenExpan.
-Unlike RetExpan which relies on intermediate semantic features of entities, GenExpan directly performs inference based on the given preceding text and the knowledge stored in the LLMs themselves, thus is more efficient.
+Unlike RetExpan which relies on intermediate semantic features of entities, GenExpan directly infers based on the given preceding text and knowledge stored in LLMs themselves, thus is more efficient.
 The proposed GenExpan consists of three phases: entity generation, entity selection, and entity re-ranking.
 In the first phase, GenExpan generates a new set of entities based on positive seed entities and current expanded entities.
-A constrained decoding strategy is employed to ensure that the new entities are included in the candidate entities.
-Subsequently, the entity selection phase filters the top-$K$ entities to be added to the current expansion, based on the similarity scores between the new entities and the positive seed entities.
-These first two steps are executed iteratively for several rounds.
-Lastly, akin to RetExpan, the final step re-ranks the previous expansion results to lower the ranking of entities identical to the negative seed entities in terms of negative attributes.
-GenExpan utilizes LLaMA-7B as the backbone.
-Additionally, we continually pre-trained LLaMA-7B to fully leverage the given corpus ${D}$.
-Figure 6. Prefix tree-based constrained decoding.
+A constrained decoding strategy is designed to ensure that the generated entities are included in the candidate entities.
+Subsequently, the entity selection phase filters $K$ entities to be added to the current expansion, based on the similarity scores between the generated entities and positive seed entities.
+The first two steps are executed iteratively for several rounds.
+Finally, akin to RetExpan, GenExpan re-ranks the previous expansion results to lower the ranking of entities identical to the negative seed entities in terms of negative attributes.
+GenExpan utilizes LLaMA-7B [21] as the backbone.
+Additionally, we continually pre-trained LLaMA-7B with the given corpus ${D}$ to fully leverage entity semantic information.
+Figure 6: Prefix tree-based constrained decoding.
 Entity Generation.
 We craft a simple prompt $\text{Prompt}_{\text{g}}$ with the given entity set to guide the model toward generating entities semantically similar to the target semantic class.
-The full prompt can be found in appendix.
+The full prompt can be found in supplementary notes.
 In the first round of generation, we randomly select 3 entities from positive seed entities $S^{pos}$ as the input entity set.
-In subsequent rounds, to maintain diversity while ensuring the semantic does not deviate from original positive seed entities, 2 entities are randomly chosen from $S^{pos}$, and 1 entity is randomly sampled from current expansion results.
-Vanilla decoding strategies like beam search (Sutskever et al., 2014) might generate entities that are not included in the candidate entities.
-Hence, we construct a prefix tree structure to ensure that the decoded entities belong to the candidate entities.
-In this structure, the root node represents the beginning, and each path from the root to a leaf node represents a complete candidate entity.
-During decoding, the process must follow a specific path from root to leaf at a time.
-In other words, for a certain node, its child nodes represent subsequent tokens that are allowed to be generated.
-By employing this prefix-constrained beam search, we ensure that the generated entities remain valid.
-Entity Selection. For each previously generated new entity $e$, we compute the positive similarity score $sco^{pos}$ between them and positive seed entities $S^{pos}$, which is calculated as follows:
-(7) $\small sco^{pos}(e)=\frac{1}{|S^{pos}|}\sum_{e^{\prime}\in S^{pos}}\sqrt[|e^{\prime}|]{P(e^{\prime}|f(e))},$
+In subsequent rounds, to maintain diversity while ensuring the semantic does not deviate from original positive seed entities, 2 entities and 1 entity are randomly sampled from $S^{pos}$ and current expansion results, respectively.
+Vanilla decoding strategies like beam search [34] might uncontrollably generate entities that are not included in the candidate entities.
+To address this issue, a prefix tree structure is constructed to constrain the decoding process.
+In this structure, the root node represents the beginning token of an entity, and each path from the root to a leaf node represents a complete candidate entity.
+During decoding, the process must follow a specific path from root to leaf at one time.
+In other words, the child nodes of each node represent the tokens allowed for subsequent generation.
+By employing this prefix-constrained beam search, we ensure the generated entities remain valid.
+Entity Selection. For each generated entity $e$, we calculate the positive similarity score $sco^{pos}$ between it and positive seed entities $S^{pos}$, which is formulated as follows:
+$sco^{pos}(e)=\frac{1}{|S^{pos}|}\sum_{e^{\prime}\in S^{pos}}\sqrt[|e^{\prime}|]{P(e^{\prime}|f(e))},$ (8)
 in which $f(e)$ is the template “{e} is similar to”.
 $P(e^{\prime}|f(e))$ represents the conditional probability of LLM generating $e^{\prime}$ when given $f(e)$.
-We use geometric mean to balance the different token numbers $|e^{\prime}|$ of various entities.
+We use geometric mean to balance the various token length $|e^{\prime}|$ of entities.
 Top-$p$ entities with the highest positive similarity scores are incorporated into the current expansion result.
 Entity Re-ranking. The process of reranking in GenExpan is identical to that in RetExpan except for the calculation of $sco^{neg}$.
-5.2.2. Strategy 1: Chain-of-thought Reasoning
-Inspired by (Wei et al., 2022), we enhance the reasoning capability of LLMs for Ultra-ESE through a series of intermediate reasoning steps.
+V-B2 Enhancement Strategy 2: Chain-of-thought Reasoning
+Inspired by [35], we enhance the reasoning capability of LLMs for Ultra-ESE through a series of intermediate reasoning steps.
 Basically, the positive seed entities $S^{pos}$ in GenExpan are directly utilized to construct $\text{Prompt}_{\text{g}}$ for generating new entities.
@@
 The chain-of-thought reasoning stimulates the reasoning capability of LLM and benefits it for perceiving the semantic classes of seed entities and the concerned attributes.
-5.2.3. Strategy 2: Entity-based Retrieval Augmentation
-Similar to RetExpan, we prepend the introduction text of the input entity before $\text{Prompt}_{\text{g}}$ to enrich the LLM’s understanding of the input entity.
-Importantly, these entity introductions are not integrated during the fine-tuning of LLaMA; they are exclusively utilized during entity generation.
-6. Experiments
-6.1. Experiment Setup
+VI Experiments
+VI-A Experiment Setup
 Compared Methods.
 We compare with three categories of models.
-The initial category encompasses traditional statistical probability-based methods, notably SetExpan (Shen et al., 2017) and CasE (Yu et al., 2019).
+The initial category encompasses traditional statistical probability-based methods, notably SetExpan [18] and CasE [36].
 The second category comprises methods based on the pre-trained language model BERT, including CGExpan and ProbExpan, with ProbExpan representing the prior state-of-the-art method.
@@
 Evaluation Metrics.
-The objective of ultra-fine-grained ESE is to expand entity list in descending order based on their similarity with positive seed entities.
-Following prior research, MAP@$K$ and P@$K$ are employed as evaluation metrics.
-MAP@$K$ is computed as follows:
+The primary objective of ultra-fine-grained ESE is to expand the entity list in descending order based on their similarity with positive seed entities.
+Following prior research, PosMAP@$K$ and PosP@$K$ are employed as evaluation metrics.
+PosMAP@$K$ is computed as follows:
 […]
-where $Q$ stands for the set of all queries, $\mbox{AP}_{K}\left(L_{q},\mathcal{P}\right)$ denotes the average precision at position $K$ with the ranked list $L_{q}$ and ground-truth list $\mathcal{P}$.
-P@$K$ is the precision of the top-$K$ entities.
+where $Q$ stands for the set of all queries, $\mbox{PosAP}_{K}\left(L_{q},\mathcal{P}\right)$ denotes the average precision at position $K$ with the ranked list $L_{q}$ and ground-truth list $\mathcal{P}$.
+PosP@$K$ is the precision of the top-$K$ entities.
 […]
-Table 2. Main experiment results. We report the MAP and P scores corresponding to Pos/Neg/Comb. The highest average score is bolded and the next highest average score is underlined.
-Method Type Method Metric Type MAP P Avg
+TABLE II: We report the MAP and P scores corresponding to Pos/Neg/Comb on UltraWiki. The highest score is bolded.
+Metric Type Method Type Method MAP P Avg
 @10 @20 @50 @100 @10 @20 @50 @100
-Probability Based SetExpan Pos $\uparrow$ 13.41 11.83 10.66 10.79 20.10 19.97 21.88 26.39 16.88
-Neg $\downarrow$ 4.06 3.77 3.71 4.20 7.66 8.10 10.92 17.44 7.48
-Comb $\uparrow$ 54.67 54.03 53.48 53.30 56.22 55.93 55.48 54.48 54.70
-CaSE Pos $\uparrow$ 16.72 13.74 11.60 10.91 24.58 23.28 26.86 30.53 19.78
-Neg $\downarrow$ 5.33 4.32 3.63 3.50 10.22 10.10 12.79 15.96 8.23
-Comb $\uparrow$ 55.69 54.71 53.99 53.70 57.18 56.59 57.03 57.28 55.77
-Retrieval Based CGExpan Pos $\uparrow$ 21.64 19.72 19.11 20.22 30.61 31.24 38.39 50.03 28.87
-Neg $\downarrow$ 6.15 6.54 8.03 9.96 12.29 16.37 27.38 41.72 16.06
-Comb $\uparrow$ 57.75 56.59 55.54 55.13 59.16 57.44 55.50 54.15 56.41
-ProbExpan Pos $\uparrow$ 21.86 22.11 22.80 23.89 38.08 39.41 47.02 62.71 34.74
-Neg $\downarrow$ 6.72 8.16 10.85 13.47 15.12 19.92 34.51 56.48 20.65
-Comb $\uparrow$ 57.57 56.98 55.97 55.21 61.48 59.75 56.25 53.12 57.04
-Generation Based GPT4 Pos $\uparrow$ 37.20 35.37 35.49 35.59 47.12 48.87 55.31 62.22 44.65
-Neg $\downarrow$ 6.04 6.61 8.03 8.35 10.40 15.06 24.57 33.63 14.09
-Comb $\uparrow$ 65.58 64.38 63.73 63.62 68.36 66.90 65.37 64.29 65.28
-Retrieval Based (Ours) RetExpan Pos $\uparrow$ 41.73 39.53 38.55 39.91 54.58 58.03 66.76 77.23 52.04
-Neg $\downarrow$ 8.77 9.04 10.65 13.29 16.44 21.04 34.78 56.54 21.32
-Comb $\uparrow$ 66.48 65.25 63.95 63.31 69.07 68.50 65.99 60.34 65.36
-+ Contrast Pos $\uparrow$ 47.45 44.68 43.63 44.20 59.83 62.02 69.36 77.92 56.14
-Neg $\downarrow$ 8.02 8.98 10.89 13.05 14.83 21.23 35.47 55.12 20.95
-Comb $\uparrow$ 69.72 67.85 66.37 65.57 72.50 70.39 66.95 61.40 67.59
-+RA Pos $\uparrow$ 44.74 42.23 40.66 40.27 58.31 61.36 66.58 73.97 53.52
-Neg $\downarrow$ 5.53 5.88 6.56 7.44 11.70 16.70 25.81 37.05 14.58
-Comb $\uparrow$ 69.60 68.18 67.05 66.42 73.31 72.33 70.39 68.46 69.47
-Generation Based (Ours) GenExpan Pos $\uparrow$ 46.79 45.00 42.89 40.80 59.77 62.15 66.26 66.57 53.78
-Neg $\downarrow$ 7.25 8.28 8.72 8.01 15.21 21.31 27.33 28.58 15.59
-Comb $\uparrow$ 69.77 68.36 67.08 66.40 72.28 70.42 69.47 68.99 69.10
-+ CoT Pos $\uparrow$ 50.39 47.80 43.67 40.06 62.74 64.45 64.06 60.38 54.19
-Neg $\downarrow$ 7.79 9.29 8.15 6.89 15.97 22.66 23.52 21.90 14.52
-Comb $\uparrow$ 71.30 69.25 67.76 66.58 73.39 70.90 70.27 69.24 69.84
-+ RA Pos $\uparrow$ 48.68 46.26 43.80 41.27 61.12 62.64 66.49 65.65 54.49
-Neg $\downarrow$ 7.16 8.54 8.63 7.70 14.79 21.18 26.67 26.84 15.19
-Comb $\uparrow$ 70.76 68.86 67.58 66.78 73.16 70.73 69.91 69.41 69.65
-6.2. Main Experiments
-The results of the main experiment are presented in Table 2, from which we can observe that:
+Pos $\uparrow$ Probability Based SetExpan 13.41 11.83 10.66 10.79 20.10 19.97 21.88 26.39 16.88
+CaSE 16.72 13.74 11.60 10.91 24.58 23.28 26.86 30.53 19.78
+Retrieval Based CGExpan 21.64 19.72 19.11 20.22 30.61 31.24 38.39 50.03 28.87
+ProbExpan 21.86 22.11 22.80 23.89 38.08 39.41 47.02 62.71 34.74
+Generation Based GPT4 37.20 35.37 35.49 35.59 47.12 48.87 55.31 62.22 44.65
+Retrieval Based (Ours) RetExpan 41.73 39.53 38.55 39.91 54.58 58.03 66.76 77.23 52.04
+RetExpan + Contrast 47.45 44.68 43.63 44.20 59.83 62.02 69.36 77.92 56.14
+Generation Based (Ours) GenExpan 46.79 45.00 42.89 40.80 59.77 62.15 66.26 66.57 53.78
+GenExpan + CoT 50.39 47.80 43.67 40.06 62.74 64.45 64.06 60.38 54.19
+Neg $\downarrow$ Probability Based SetExpan 4.06 3.77 3.71 4.20 7.66 8.10 10.92 17.44 7.48
+CaSE 5.33 4.32 3.63 3.50 10.22 10.10 12.79 15.96 8.23
+Retrieval Based CGExpan 6.15 6.54 8.03 9.96 12.29 16.37 27.38 41.72 16.06
+ProbExpan 6.72 8.16 10.85 13.47 15.12 19.92 34.51 56.48 20.65
+Generation Based
+GPT4 6.04 6.61 8.03 8.35 10.40 15.06 24.57 33.63 14.09
+Retrieval Based (Ours) RetExpan 8.77 9.04 10.65 13.29 16.44 21.04 34.78 56.54 21.32
+RetExpan + Contrast 8.02 8.98 10.89 13.05 14.83 21.23 35.47 55.12 20.95
+Generation Based (Ours) GenExpan 7.25 8.28 8.72 8.01 15.21 21.31 27.33 28.58 15.59
+GenExpan + CoT 7.79 9.29 8.15 6.89 15.97 22.66 23.52 21.90 14.52
+Comb $\uparrow$ Probability Based SetExpan 54.67 54.03 53.48 53.30 56.22 55.93 55.48 54.48 54.70
+CaSE 55.69 54.71 53.99 53.70 57.18 56.59 57.03 57.28 55.77
+Retrieval Based CGExpan 57.75 56.59 55.54 55.13 59.16 57.44 55.50 54.15 56.41
+ProbExpan 57.57 56.98 55.97 55.21 61.48 59.75 56.25 53.12 57.04
+Generation Based
+GPT4 65.58 64.38 63.73 63.62 68.36 66.90 65.37 64.29 65.28
+Retrieval Based (Ours) RetExpan 66.48 65.25 63.95 63.31 69.07 68.50 65.99 60.34 65.36
+RetExpan + Contrast 69.72 67.85 66.37 65.57 72.50 70.39 66.95 61.40 67.59
+Generation Based (Ours) GenExpan 69.77 68.36 67.08 66.40 72.28 70.42 69.47 68.99 69.10
+GenExpan + CoT 71.30 69.25 67.76 66.58 73.39 70.90 70.27 69.24 69.84
+VI-B Main Experiments
+The results of the main experiment are presented in Table II, from which we can observe that:
 (1) Regarding the average values of the comprehensive metrics CombMAP and CombP, our proposed GenExpan achieves optimal and sub-optimal performances with the incorporation of the chain-of-thought reasoning strategy and the entity-based retrieval augmentation strategy, respectively.
@@
 This insight prompts future research into devising better representations to express the ultra-fine-grained semantics of entities, such as decoupling the base semantics of entities from the ultra-fine-grained attribute semantics, similar to the Mix-of-Expert (MoE) approach, where distinct features represent different perspectives of the semantics.
-The ablation experiments presented in Table 3 further validate our hypothesis regarding the factors why RetExpan and GenExpan surpass the previous state-of-the-art models.
+The ablation experiments presented in Table III further validate our hypothesis regarding the factors why RetExpan and GenExpan surpass the previous state-of-the-art models.
 Both methods exhibit remarkable degradation after removing the corresponding modules.
-(3) The simple retrieval augmentation strategy yielded remarkable results for both RetExpan and GenExpan.
-Particularly for RetExpan, the strategy improves the comprehensive Comb metrics by 4.11 points on average.
-The strategy primarily optimizes the Neg class metrics, leading to a significant reduction in the intrusion of negative target entities.
-However, its effect on Pos class metrics exhibits instability.
-For instance, the PosMAP@50 and PosMAP@100 of RetExpan+RA are instead lower compared to the original RetExpan.
-This phenomenon can be attributed to the fact that the extra introductions corresponding to each entity are currently fixed and do not adapt to changes in the entity context.
-Consequently, this increases the homogeneity of the entity embedding between irrelevant entities (neither positive nor negative target entities) and positive seed entities under the same fine-grained semantic class to some extent.
-Our point is further supported by the observation that RetExpan’s MAP@100 for fine-grained semantic classes grows from 82.08 to 87.17 after the introduction of retrieval augmentation strategy.
-This suggests that a dynamic and more fine-grained knowledge retrieval strategy needs further investigation.
-(4) Among all the enhancement strategies, ultra-fine-grained contrastive learning provides the most substantial improvement in the Pos class metrics, averaging 4.10 points (from 52.04 to 56.14).
+(3) Comparing the two enhancement strategies, ultra-fine-grained contrastive learning provides the most substantial improvement in the Pos class metrics, averaging 4.10 points (from 52.04 to 56.14).
 Conversely, contrastive learning brings smaller gains in Neg class metrics.
@@
 However, this does not imply that these methods exhibit a high degree of negative semantic awareness.
-The Pos class metrics of SetExpan and CaSE are similarly low, indicating their limited understanding of fine-grained semantic classes overall.
+The Pos class metrics of SetExpan and CaSE are low in similarity, indicating their limited understanding of fine-grained semantic classes overall.
 Consequently, they struggle to recall entities with the given seed entities, leading to low scores in both Pos and Neg class metrics.
@@
 (5) Without further fine-tuning and modifications to the model structure, GPT-4 achieves excellent results, which proves that generative LLMs have great potential for addressing Ultra-ESE tasks.
-However, GPT-4 still fails to outperform the GenExpan, which is also generative and based on smaller LLaMA-7B.
+However, GPT-4 still fails to outperform the GenExpan, which is also generative but based on smaller LLaMA-7B.
 Analysis of cases reveals that there are two primary problems with GPT-4.
 On the one hand, it performs poorly on long-tail problems, such as U.S.
-national monuments, mobile phone brands, which contain a considerable number of low-frequency entities with limited information available on the Internet.GPT-4 only achieves single-digit PosMAP on these semantic classes.
-In contrast, GenExpan performs better, benefiting from the given contextual corpus and targeted retrieval augmentation strategies.
-On the other hand, GPT-4 is prone to haphazardly generate non-existent entities (e.g., fake mobile phone brands), which is referred to hallucination problem in recent work(Zhang et al., 2023a; Ji et al., 2023).
+national monuments and mobile phone brands, which contain a considerable number of low-frequency entities with limited information available on the Internet.
+GPT-4 only achieves single-digit PosMAP on these semantic classes.
+In contrast, GenExpan performs better, benefiting from the given contextual corpus.
+On the other hand, GPT-4 is prone to haphazardly generate non-existent entities (e.g., fake mobile phone brands), which is referred to hallucination problem in recent work[37, 38].
 We are currently unable to solve this issue by simple output post-processing.
 While the results of GPT-4 are impressive, there remains a lot of space for improvement.
-Table 3. Ablation experiments for each module of RetExpan and GenExpan. Each module contributes to the overall performance.
+TABLE III: Ablation experiments for each module of RetExpan and GenExpan. Each module contributes to the overall performance.
 Method MAP Avg
@@
 - Further pretrain 68.58 66.67 65.23 64.23 66.18
-6.3. Analysis of Negative Entities
+VI-C Analysis of Negative Entities
 We conducted analytical experiments from various perspectives regarding negative seed entities to answer the following questions:
@@
 Thanks to the high scalability, it was also integrated into ProbExpan.
-Results of the three comparison experiments are presented in Table 5.
+Results of the three comparison experiments are presented in Table IV.
 We can observe that:
@@
 This strongly suggests that the introduction of negative seed entities effectively mitigates the intrusion of negative target entities while raising the ranking of positive target entities, i.e., portraying the semantic classes at a finer granularity.
-(2) Further analyzing the Pos and Neg metrics, we observe that the addition of the entity re-ranking module to ProbExpan conversely resulted in a notable decline in the PosP metrics.
-It’s essential to clarify the distinction between the MAP and P metrics: P@K is concerned solely with the number of target entities in the top-K entity list, regardless of their positions, while MAP@K is rank-aware, with higher values indicating that target entities are positioned closer to the top of the list.
+(2) Further analyzing the Pos and Neg metrics, we observe that the addition of entity re-ranking module to ProbExpan conversely resulted in a notable decline in PosP metrics.
+It’s essential to clarify the distinction between MAP and P metrics: P@K is concerned solely with number of target entities in the top-K entity list, regardless of their positions, while MAP@K is rank-aware, with higher values indicating that target entities are positioned closer to the top of the list.
 Consequently, this indicates that the incorporation of negative seed entities in ProbExpan makes a mixed impact, propelling some positive target entities to higher ranks and simultaneously causing others to be ranked lower.
 (3) For RetExpan, negative seed entities have a similar impact on metrics with different K values, whereas for GenExpan, negative seed entities have a greater impact on metrics with a smaller K.
-This may be due to the fact that RetExpan is a one-time expansion (Kushilevitz et al., 2020; Mamou et al., 2018; Yu et al., 2019) model and GenExpan is an iterative expansion (Huang et al., 2020; Rong et al., 2016; Shen et al., 2017) model, so entities expanded later (corresponding to a larger K) are more likely to deviate from the original ground-truth semantic class.
+This may be due to the fact that RetExpan is a one-time expansion [39, 40, 36] model and GenExpan is an iterative expansion [41, 42, 18] model, so entities expanded later (corresponding to a larger K) are more likely to deviate from the original ground-truth semantic class.
 During these stages, re-ranking using negative seed entities also doesn’t work.
-Addressing the issue of semantic drift (Curran et al., 2007; McIntosh, 2010; Shi et al., 2014) in iterative expansion models like GenExpan is a longstanding challenge in ESE field.
-Table 4. Comparison experiments when positive and negative attributes are the same and different.
-Method Metric Type MAP Avg
-@10 @20 @50 @100
-$\mathbf{\mathcal{A}^{pos}=\mathcal{A}^{neg}}$
-RetExpan Pos 43.14 41.74 41.93 42.89 42.43
-Neg 5.54 6.44 8.71 11.56 8.06
-Comb 68.80 67.65 66.61 65.67 67.18
-RetExpan + Contrast Pos 49.68 48.73 48.77 49.50 49.17
-Neg 5.76 6.51 8.64 11.62 8.13
-Comb 71.96 71.11 70.06 68.94 70.52
-RetExpan + RA Pos 46.29 44.03 43.46 43.61 44.35
-Neg 5.45 5.86 6.84 7.56 6.43
-Comb 70.42 69.08 68.31 68.03 68.96
-$\mathbf{\mathcal{A}^{pos}\neq\mathcal{A}^{neg}}$
-RetExpan Pos 41.73 39.53 38.55 39.91 39.93
-Neg 8.77 9.04 10.65 13.29 10.44
-Comb 66.48 65.25 63.95 63.31 64.75
-RetExpan + Contrast Pos 46.44 43.69 42.51 43.14 43.95
-Neg 8.22 9.25 11.17 13.38 10.51
-Comb 69.11 67.22 65.67 64.88 66.72
-RetExpan + RA Pos 44.74 42.23 40.66 40.27 41.98
-Neg 5.53 5.88 6.56 7.44 6.35
-Comb 69.60 68.17 67.05 66.42 67.81
-Table 5. Ablation experiments on entity re-ranking module with negative seed entities.
+Addressing the issue of semantic drift [25, 43, 27] in iterative expansion models like GenExpan is a longstanding challenge in the ESE field.
+TABLE IV: Ablation experiments on entity re-ranking module with negative seed entities.
 MAP P
@@
 $\triangle$ Comb -0.65 -0.56 -0.59 -0.57 -0.50 0.00 -0.10 0.00 -0.37
-What are the roles played by negative seed entities?
-We first assessed the efficacy of RetExpan and its enhancement strategies by comparing scenarios where positive and negative attributes are identical or different.
-As shown in Table 4, we can find that:
-(1) When positive and negative attributes are identical, there is no overlap between positive and negative target entities, which simplifies semantic classes and results in relatively high performance.
-In such cases, negative seed entities play a supportive role, primarily emphasizing user-concerned attributes and mitigating ambiguity.
-Conversely, when positive and negative attributes are different, the potential overlap between entities satisfying these attributes poses a significant challenge to semantic understanding, leading to lower overall performance.
-Here, negative seed entities play a crucial role in conveying unwanted semantics and are indispensable.
-(2) The benefits derived from contrastive learning are more pronounced in semantic classes where positive and negative attributes are identical.
-This is because, in such scenarios, there is no overlap between positive and negative target entities, making the contrastive training pairs mined based on the similarity of positive and negative seed entities more reliable.
-(3) When positive and negative attributes differ, the primary focus lies in evaluating the model’s capability to comprehend the semantics expressed by negative seed entities.
-We observe that retrieval augmentation demonstrates more effective in this situation.
-This phenomenon suggests that the understanding of negative seed entities demands more additional knowledge compared to positive seed classes.
-This notion is further verified in the analysis of chain-of-thought reasoning in Section 6.4.
-What are the differences in challenges posed by semantic classes with varying numbers of positive and negative attributes?
-We analyze the performance of RetExpan on semantic classes characterized by varying quantities of positive and negative attributes in Table 6.
-To maintain control over variables, we exclusively present semantic classes for three combinations of attribute quantity: (1,1), (1,2), and (2,1).
-Imposing stricter constraints on the attributes results in a reduction in the number of associated target entities, rendering it more challenging to discern these entities accurately.
-Consequently, a decrease in the Pos (Neg) metrics is observed, correlating with an increase in the number of positive (negative) attributes.
-It’s important to emphasize that although the Neg metrics decrease with more negative attribute constraints, this does not imply a reduced difficulty in distinguishing negative target entities.
-Rather, identifying these negative target entities requires the model to possess more comprehensive attribute information about the negative seed entities.
-Table 6. Comparison experiments on semantic classes with different numbers of positive and negative attributes.
+TABLE V: Comparison experiments when positive and negative attributes are the same and different.
+Method Metric Type MAP Avg
+@10 @20 @50 @100
+$\mathbf{\mathcal{A}^{pos}=\mathcal{A}^{neg}}$
+RetExpan Pos 43.14 41.74 41.93 42.89 42.43
+Neg 5.54 6.44 8.71 11.56 8.06
+Comb 68.80 67.65 66.61 65.67 67.18
+RetExpan + Contrast Pos 49.68 48.73 48.77 49.50 49.17
+Neg 5.76 6.51 8.64 11.62 8.13
+Comb 71.96 71.11 70.06 68.94 70.52
+$\mathbf{\mathcal{A}^{pos}\neq\mathcal{A}^{neg}}$
+RetExpan Pos 41.73 39.53 38.55 39.91 39.93
+Neg 8.77 9.04 10.65 13.29 10.44
+Comb 66.48 65.25 63.95 63.31 64.75
+RetExpan + Contrast Pos 46.44 43.69 42.51 43.14 43.95
+Neg 8.22 9.25 11.17 13.38 10.51
+Comb 69.11 67.22 65.67 64.88 66.72
+TABLE VI: Comparison experiments on semantic classes with different numbers of positive and negative attributes.
 $(|\mathcal{A}^{pos}|,|\mathcal{A}^{neg}|)$ Metric Type MAP Avg
@@
 Comb 68.22 64.78 65.51 65.76 66.07
-6.4. Analysis of Enhancement Strategy
-Table 7. Ablation experiments on contrastive learning.
+What are the roles played by negative seed entities?
+We first assessed the efficacy of RetExpan and its enhancement strategies by comparing scenarios where positive and negative attributes are identical or different.
+As shown in Table V, we can find that:
+(1) When positive and negative attributes are identical, there is no overlap between positive and negative target entities, which simplifies semantic classes and yielding higher performance.
+In such cases, negative seed entities play a supportive role, primarily emphasizing user-concerned attributes and mitigating ambiguity.
+Conversely, when positive and negative attributes are different, the potential overlap between entities satisfying these attributes poses a significant challenge to semantic understanding, leading to lower overall performance.
+Here, negative seed entities play a crucial role in conveying unwanted semantics and are indispensable.
+(2) The benefits derived from contrastive learning are more pronounced in semantic classes where positive and negative attributes are identical.
+This is because, in such scenarios, there is no overlap between positive and negative target entities, making the contrastive training pairs mined based on the similarity of positive and negative seed entities more reliable.
+(3) When positive and negative attributes differ, the primary focus lies in evaluating the model’s capability to comprehend the semantics expressed by negative seed entities.
+We observe that contrastive learning demonstrates less effective in this situation.
+This phenomenon suggests that the understanding of negative seed entities demands more additional knowledge compared to positive seed classes.
+This notion is further verified in the analysis of chain-of-thought reasoning in Section VI-D.
+What are the differences in challenges posed by semantic classes with varying numbers of positive and negative attributes?
+We analyze the performance of RetExpan on semantic classes characterized by varying quantities of positive and negative attributes in Table VI.
+To maintain control over variables, we exclusively present semantic classes for three combinations of attribute quantity: (1,1), (1,2), and (2,1).
+Imposing stricter constraints on the attributes results in a reduction in the number of associated target entities, rendering it more challenging to discern these entities accurately.
+Consequently, a decrease in the Pos (Neg) metrics is observed, correlating with an increase in the number of positive (negative) attributes.
+It’s important to emphasize that although the Neg metrics decrease with more negative attribute constraints, this does not imply a reduced difficulty in distinguishing negative target entities.
+Rather, identifying these negative target entities requires more comprehensive attribute information about negative seed entities.
+VI-D Analysis of Enhancement Strategy
+TABLE VII: Ablation experiments on contrastive learning.
 Method Metric Type MAP Avg
@@
 Comb 69.211 67.221 65.8875 65.0875 66.85
-6.4.1. Analysis of Contrastive Learning
+VI-D1 Analysis of Contrastive Learning
 We conducted ablation experiments on each part of the training data for ultra-fine-grained contrastive learning to explore their impact on overall performance.
-As illustrated in Table 7, based on the full training data, the last three rows sequentially remove the hard negative samples comprising pairs of positive and negative target entities, the normal negative samples consisting of pairs of positive and negative target entities along with entities from other semantic classes, and the positive samples composed of entity pairs within the same ultra-fine-grained semantic class.
+As illustrated in Table VII, based on the full training data, the last three rows sequentially remove the hard negative samples comprising pairs of positive and negative target entities, the normal negative samples consisting of pairs of positive and negative target entities along with entities from other semantic classes, and the positive samples composed of entity pairs within the same ultra-fine-grained semantic class.
 It can be seen that hard negative samples consisting of pairs of positive and negative target entities contribute the most to the overall performance, which is in line with our intuition of proposing ultra-fine-grained contrastive learning.
-We set both —$L_{pos}$— and —$L_{neg}$— to 10 to mitigate the introduction of excessive noise.
+We set both $|L_{pos}|$ and $|L_{neg}|$ to 10 to mitigate the introduction of excessive noise.
 Consequently, it’s evident that this part of training data wields more influence when K is small.
-Exploring how to mitigate the noise introduced by automatic GPT-4 labeling when —$L_{pos}$— and —$L_{neg}$— increase is a promising direction for future exploration.
+Exploring how to mitigate the noise introduced by automatic GPT-4 labeling when $|L_{pos}|$ and $|L_{neg}|$ increase is a promising direction for future exploration.
 Additionally, we observe that normal negative samples boost the overall performance.
@@
 Moreover, positive samples, composed of entities belonging to the same ultra-fine-grained semantic class, similarly contribute positively to the overall performance.
-Table 8. Comparison experiments with different retrieval contents on retrieval augmentation.
-Method Metric Type MAP Avg
-@10 @20 @50 @100
-RetExpan + RA (Entity Introduction) Pos 44.74 42.23 40.66 40.27 41.97
-Neg 5.53 5.88 6.56 7.44 6.35
-Comb 69.60 68.18 67.05 66.42 67.81
-RetExpan + RA (Wikidata Attributes) Pos 40.94 38.78 37.28 37.64 38.66
-Neg 6.93 6.79 7.89 9.56 7.79
-Comb 67.00 66.00 64.69 64.04 65.43
-RetExpan + RA (GT Attributes) Pos 45.28 43.84 43.77 45.22 44.52
-Neg 6.21 6.69 7.92 8.19 7.25
-Comb 69.54 68.57 67.92 68.51 68.64
-GenExpan + RA (Entity Introduction) Pos 48.68 46.26 43.80 41.27 45.00
-Neg 7.16 8.54 8.63 7.70 8.01
-Comb 70.76 68.86 67.58 66.78 68.50
-GenExpan + RA (Wikidata Attributes) Pos 48.30 45.94 43.91 41.50 44.91
-Neg 7.00 8.46 8.79 7.83 8.02
-Comb 70.65 68.74 67.56 66.83 68.45
-GenExpan + RA (GT Attributes) Pos 49.14 46.83 43.98 42.51 45.62
-Neg 6.76 8.43 8.56 7.66 7.85
-Comb 71.19 69.20 67.71 67.42 68.88
-6.4.2. Analysis of Retrieval Augmentation
-In this subsection, we delve into the impact of additional entity knowledge retrieved from various sources on RetExpan and GenExpan, including entity introductions and attribute information extracted from Wikidata, as well as ground-truth positive and negative attribute information.
-The experiment results presented in Table 8 demonstrate that each of the three knowledge is beneficial for both the original RetExpan and GenExpan.
-Concerning the entity attribute information sourced from Wikidata, we randomly sampled an attribute subset and converted it into text with templates.
-The attributes on Wikidata are intricate and encompass numerous rarely-used attributes (e.g., the attributes of the entity ”Motorola” include its ”YouTube channel ID”).
-This may lead to the retrieved attributes are not positive or negative attributes that the current ultra-fine-grained semantic class really cares about, thereby hindering model comprehension.
-Therefore, despite the high quality of Wikidata’s attribute information, its gain to the model is relatively marginal.
-Meanwhile, we can easily find that the direct inclusion of ground-truth positive and negative attribute information related to ultra-fine-grained semantic classes has the most substantial gain to the model.
-Yet, we posit that the potential of retrieval augmentation methods has not been fully realized.
-Firstly, the supplementary knowledge retrieved for each entity is static across different sentences and does not adapt to the entity’s context, potentially leading to homogenized entity representations.
-Crafting dynamic and ultra-fine-grained retrieval strategies deserves further exploration.
-Secondly, simply concatenating retrieved knowledge with the original sentences is not the optimal way to leverage it.
-Exploring more effective methods for knowledge exploitation is also essential.
-6.4.3. Analysis of Chain-of-thought Reasoning
-Table 9.
-Comparison experiments on chain-of-thought reasoning.
+VI-D2 Analysis of Chain-of-thought Reasoning
+TABLE VIII: Comparison experiments on chain-of-thought reasoning.
 The impact of reasoning depth and reasoning precision on performance is investigated.
@@
 Comb 70.98 69.00 67.38 66.46 68.46
-GenExpan + CoT (Cen CN + Gen Pos) Pos 50.39 47.80 43.67 40.06 45.48
+GenExpan + CoT (Cen CN & Gen Pos) Pos 50.39 47.80 43.67 40.06 45.48
 Neg 7.79 9.29 8.15 6.89 8.03
 Comb 71.30 69.25 67.76 66.58 68.72
-GenExpan + CoT (Cen CN + GT Pos) Pos 50.77 48.25 44.38 40.67 46.02
+GenExpan + CoT (Cen CN & GT Pos) Pos 50.77 48.25 44.38 40.67 46.02
 Neg 7.68 9.16 8.12 6.83 7.95
 Comb 71.55 69.54 68.13 66.92 69.04
-GenExpan + CoT (Cen CN + Gen Pos + Gen Neg) Pos 49.78 47.24 42.68 39.01 44.68
+GenExpan + CoT (Cen CN & Gen Pos & Gen Neg) Pos 49.78 47.24 42.68 39.01 44.68
 Neg 7.60 8.97 7.98 6.75 7.82
 Comb 71.09 69.14 67.35 66.13 68.43
-GenExpan + CoT (Cen CN + GT Pos + GT Neg) Pos 51.02 49.19 44.86 41.40 46.62
+GenExpan + CoT (Cen CN & GT Pos & GT Neg) Pos 51.02 49.19 44.86 41.40 46.62
 Neg 7.30 8.70 7.59 6.53 7.53
 Comb 71.86 70.25 68.63 67.43 69.54
-Experiments in Table 9 examine the chain-of-thought with varying reasoning depths and assess the impact of reasoning precision on model performance. From the results, we observe that:
-(1) Deeper reasoning proves to be more advantageous for Ultra-ESE, as it facilitates a more explicit understanding of the implicit ultra-fine-grained semantics of seed entities.
+Experiments in Table VIII examine the chain-of-thought with varying reasoning depths and assess the impact of reasoning precision on model performance.
+From the results, we observe that: (1) Deeper reasoning proves to be more advantageous for Ultra-ESE, as it facilitates a more explicit understanding of the implicit ultra-fine-grained semantics of seed entities.
 (2) Surprisingly, the use of manually labeled ground-truth class names is not superior to those derived through reasoning with the LLaMA-7B.
-Case study of the generated class names reveal that LLaMA-7B can produce class names encapsulating positive attribute information based on the given positive seed entities.
+The case study of the generated class names reveals that LLaMA-7B can produce class names encapsulating positive attribute information based on the given positive seed entities.
 For instance, when given entities about U.S.
@@
 This difficulty is particularly pronounced for long-tailed entities (such as ancient musical instruments) and semantic classes with diverse attributes, which indicates that enhancing the ultra-fine-grained perception of entities within LLMs is a promising avenue for future research.
-6.5. Explore Experiments on Different Paradigm Model Interactions
-Table 10. Experiments on the interaction of RetExpan and GenExpan.
-Method Metric Type MAP Avg
-@10 @20 @50 @100
-RetExpan Pos 41.73 39.53 38.55 39.91 39.93
-Neg 8.77 9.04 10.65 13.29 10.44
-Comb 66.48 65.25 63.95 63.31 64.75
-RetExpan + GenExpan Pos 45.93 44.37 42.96 42.35 43.90
-Neg 7.03 7.51 9.04 9.48 8.27
-Comb 69.45 68.43 66.96 66.43 67.82
-GenExpan Pos 46.84 44.99 42.88 40.79 43.88
-Neg 7.27 8.29 8.74 8.02 8.08
-Comb 69.79 68.35 67.07 66.38 67.90
-GenExpan + RetExpan Pos 48.91 46.09 44.13 42.49 45.41
-Neg 6.24 6.80 7.56 7.84 7.11
-Comb 71.33 69.64 68.29 67.32 69.15
-We are curious whether retrieval-based and generation-based frameworks can interact and mutually reinforce each other.
-Actually, in the process of enhancing the retrieval-based RetExpan, we have leveraged the generative LLMs GPT-4.
-At the same time, when enhancing the generation-based GenExpan, we have utilized the retrieval-based approach to retrieve extra entity knowledge.
-In this section, we explore the facilitative relationship between the two paradigms at a higher level.
-Specifically, for the A+B model in Table 10, we first allow Model A to obtain 1000 entities from the complete candidate entities, which guarantees high entity recall.
-Then, model B executes on the obtained 1000 entities.
-To our surprise, no matter whether RetExpan or GenExpan is deployed first, both of them enhance the performance on all metrics considerably.
-These two different paradigm frameworks have their own advantages.
-The retrieval-based framework measures the semantic relation between candidate and seed entities by feature similarity, which inherently fits the nature of the retrieval task.
-Thus only by utilizing smaller models, retrieval-based frameworks can achieve competitive results, and be easily migrated to specific domains.
-Generation-based frameworks, on the other hand, typically rely on LLMs, which offer a broader knowledge base and greater semantic awareness, but the LLMs tend to be less efficient.
-Therefore, how to make the two types of models learn collaboratively, promote each other, and then achieve automatic evolution is a promising direction for solving Ultra-ESE.
-Appendix A Prompts
-When constructing contrastive learning training data, we use GPT’s classification to obtain $L_{pos}$ and $L_{neg}$.
-For each entity in the top-$T$ of $L_{0}$, determine whether its attributes are consistent with $S^{pos}$ ($S^{neg}$).
-If it is consistent, GPT should output 1; otherwise, GPT should output 0.
-Next, these entities that GPT deems consistent with the attributes of $S^{pos}$ ($S^{neg}$) will be merged with $S^{pos}$ ($S^{neg}$) to form $L_{pos}$ ($L_{neg}$).
-The prompt to classify entity is shown in Table 13.
-In GenExpan, each round uses three example entities to construct $Prompt_{g}$, which is used to generate entities that are semantically similar to them. $Prompt_{g}$ is as shown in Table 14.
-In the process of enhancing GenExpan with Chain of Thought prompting, we employ LLM to generate fine-grained class names. The prompt, $Prompt_{c}$, is shown in Table 15.
-Appendix B Implementation Details of RetExpan
-In RetExpan, sentences are tokenized using the WordPiece tokenizer and then fed into a 12-layer Transformer initialized with $BERT_{BASE}$ weights.
-To optimize training efficiency and preserve semantic knowledge learned by BERT, we freeze the first 11 layers of the encode, thus only the last layer is fine-tuned.
-When tokenizing, to ensure that mentions of entities in long sentences are not truncated, we have implemented entity focus, guaranteeing the presence of entities within the sentence.
-During expansion, we use one-shot expansion to directly obtain the preliminary expansion results $L_{0}$.
-In RetExpan with contrastive learning strategy, each epoch during training alternates between computing and optimizing entity prediction loss and contrastive learning loss.
-To ensure that the encoder acquires knowledge from the corpus, we trained it for 20 epochs on 8 RTX 3090 GPUs.
-The hyperparameters for training learning rate, batch size, weight decay, label smoothing factor $\eta$ were set to 4e-5, 128, 1e-2, and 0.075 respectively.
-Appendix C Implementation Details of GenExpan
-In GenExpan, we use LLaMA-7b as the base model.
-Initially, we train for 1 epoch on the corpus using 6 A100 GPUs.
-The training hyperparameters learning rate, batch size, gradient accumulation steps, weight decay, gradient clipping are set to 1e-5, 4, 8, 1e-4, and 1.0 respectively.
-During entity generation in GenExpan, we utilize prefix-constrained beam search with a beam size of 40 to generate 40 entities at a round.
-For entity selection, we compute the positive similarity score for each entity and select those whose scores are in the top 0.7 as the results for the current round.
-When no new entity is generated for 20 consecutive rounds, the generation process will end and move to re-ranking.
-Appendix D Details of UltraWiki
-UltraWiki contains 10 fine-grained semantic classes, which fall under 5 coarse-grained semantic categories: Organization, Location, Product, Person, and Miscellaneous.
-The number of entities in each fine-grained semantic class ranges from 45 to 952.
-Each fine-grained semantic class has 2 to 3 attributes, and depending on the combination of attributes, several ultra-fine-grained semantic classes can be derived from a single fine-grained class.
-Detailed data is displayed in Table 11.
-The number of combinations from positive and negative attributes varies significantly, leading to substantial differences in the number of ultra-fine-grained semantic classes.
-The count of UltraWiki’s positive and negative attributes and their corresponding ultra-fine-grained classes are presented in Table 12.
-Figure 7. Parameter analysis experiments.
-Table 11. Fine-grained semantic classes detail. CLS.: Semantic Class
-Coarse CLS. Fine-grained CLS. #Entities #Ultra-fine-grained CLS. Attributes
-Organization Canada universities 99 10 ¡Loc-Province¿, ¡Type¿
-Location China cities 675 50 ¡Province¿, ¡Prefecture¿
-Countries 190 68 ¡Continent¿, ¡Driving-Side¿, ¡Per-Capita-Income¿
-US airports 370 74 ¡Role¿, ¡Loc-State¿
-US national monuments 112 12 ¡Loc-State¿ ¡Agency¿
-Product Mobile phone brands 159 7 ¡Loc-Continent¿, ¡Status¿
-Percussion instruments 128 10 ¡Type¿, ¡Source-Continent¿
-Person Nobel laureates 952 11 ¡Prize¿, ¡Gender¿
-US presidents 45 5 ¡Party¿, ¡Birth-State¿
-Miscellaneous Chemical elements 118 14 ¡Period¿, ¡Phase-at-R.T.¿
-Table 12. Types of ultra-fine-grained semantic classes. CLS.: Semantic Class
-$|\mathcal{A}^{pos}|$ $|\mathcal{A}^{neg}|$Ultra-fine-grained CLS.
-1 1 238
-1 2 5
-2 1 9
-2 2 7
-3 3 2
-Appendix E Parameter analysis
-To explore the sensitivity of each key hyperparameter in RetExpan and GenExpan, we conducted a comprehensive analysis of the smoothing factor $\eta$ and the re-ranking segment length $l$ in RetExpan, as well as the number of positive and negative entities mined in the contrastive learning strategy $|L_{pos}|$ and $|L_{neg}|$, the Top-$p$ of entity selection module in GenExpan, and the re-ranking segment length $l$, respectively.
-The results are depicted in Figure 7, from which we can observe that:
-(1) The label smoothing $\eta$ serves to mitigate the penalty for entities possessing similar semantics to ground-truth entities.
-In Ultra-ESE, there are fewer entities exhibiting similarity to ground-truth entities at the ultra-fine-grained semantic level, thereby diminishing the utility of label smoothing.
-Consequently, the fluctuation in model performance due to parameter variations is relatively not significant.
-(2) Excessive segment length $l$ in both RetExpan and GenExpan equates to degrading the segmented re-ranking strategy to a naive re-ranking approach, thereby introducing noise from irrelevant entities, as mentioned in the main text.
-Hence, there is a general decrease in both PosMAP@K as the segment length increases, with larger K values leading to greater declines.
-(3) Regarding the ultra-fine-grained contrastive learning strategy in RetExpan, we observe that mining too many or too few positive and negative entities using GPT-4 results in performance degradation.
-Insufficient positive and negative entities fail to fully stimulate the potential of contrast learning, whereas an excess of such entities introduces noisy entities.
-Therefore, as the number of entities to be mined increases gradually, PosMAP initially rises and then falls, and NegMAP exhibits a slight initial decline followed by a rise.
-However, overall, these two parameters demonstrate robustness, with even the worst-performing set of values outperforming the original RetExpan.
-(4) A trade-off is also necessary for the Top-$p$.
-A too small Top-$p$ may fail to ensure the diversity of newly generated entities, while a too large Top-$p$ may introduce an excessive number of irrelevant entities during iterative expansion.
-Consequently, we observe a decrease in PosMAP and a notable increase in NegMAP when Top-p is increased from 80% to 90%, which indicates that there is considerable space for improving the current rejection strategy for negatively targeted entities.
-Appendix F Analysis of Model Size
-We utilize various families and sizes of LLMs as backbone model of GenExpan, including 560M, 1B7, 3B, 7B1 for BLOOM, as well as 7B and 13B for LLaMA.
-The experiment results depicted in Figure 8 align well with the expectation.
-Both the BLOOM and LLaMA families satisfy the scaling law: larger models tend to be more effective in Ultra-ESE, yielding better results.
-Moreover, LLaMA-7B outperforms BLOOM-7B1 at the same scale.
-Figure 8. Comparison experiments on different LLM families and sizes.
-Figure 9.
-Case studies of GenExpan and the retrieval augmentation, chain-of-thought reasoning techniques.
+Figure 7: Case studies of GenExpan and the chain-of-thought reasoning technique.
 $+$ and - - - stand for positive and negative target entities, respectively.
@@
 represents irrelevant entities belonging to the same fine-grained semantic class as the seed entities.
-Appendix G Case Study
-Figure 9 presents some case studies of GenExpan and the corresponding retrieval augmentation, chain-of-thought reasoning techniques.
+VI-E Case Study
+Figure 7 presents some case studies of GenExpan and the chain-of-thought reasoning technique.
 Overall, it is evident that GenExpan predominantly avoids expanding entities that do not belong to the same semantic class as the positive and negative seed entities.
 This emphasizes the relative tractability of traditional ESE tasks compared to the significant challenges posed by Ultra-ESE for existing LLMs, where the model cannot exclude the expansion of negative target entities and irrelevant entities (belonging to the same fine-grained semantic class).
-For retrieval augmentation, more positive target entities are introduced after fetching the additional knowledge.
-However, for some negative target entities that fulfill positive attributes, the retrieved knowledge may contain the introduction about its positive attributes, conversely misleading the model to overlook its negative attributes.
-For example, for the negative target entity “Xinxiang” in the second column, it also belongs to Henan province, i.e., it satisfies the requirement of positive attributes.
-Therefore, after concatenating the retrieved entity introduction “Xinxiang is a prefecture-level city in northern Henan province, China.”, it is incorrectly introduced instead.
 For chain-of-thought reasoning, an interesting example is shown in Figure 7.
@@
 This also shows the potential for improvement in the chain-of-thought reasoning strategy.
-Table 13. The prompt used to select the entities that are consistent with the attributes of $S^{pos}$ ($S^{neg}$) in the top-$T$ entities of $L_{0}$ to construct $L_{pos}$ ($L_{neg}$).
-I have a task that involves classifying candidate entities based on their alignment with a seed entity set.
-The seed entities are grouped together because they share certain attributes, referred to as seed attributes.
-I will provide a list of seed entities along with their seed attributes.
-Additionally, I have a list of candidate entities that are similar to the seed entities but may not necessarily share the same seed attributes.
-I need you to identify the seed attributes and use them to classify each candidate entity into one of two categories: 1) consistent with the seed entity set in terms of seed attributes, or 0) inconsistent with the seed entity set in terms of seed attributes.
-For the given N candidate entities, please output N values, each being 1 or 0, indicating whether each candidate is consistent (1) or inconsistent (0) with the seed entity set based on the seed attributes.
-Input:
-Seed entities: [Mark Twain, Ernest Hemingway, F. Scott Fitzgerald]
-Candidate entities: [J.K. Rowling, Stephen King, Agatha Christie, John Steinbeck, Harper Lee, Charles Dickens, Virginia Woolf], total 7 entities
-Output:
-”result”: [0,1,0,1,1,0,0]
-Input:
-Seed entities: [Golden Retriever, German Shepherd, Labrador Retriever]
-Candidate entities: [Bengal Tiger, Beagle, Siberian Husky, African Elephant, Pug], total 5 entities
-Output:
-”result”: [0,1,1,0,1]
-Input:
-Seed entities: [{Entity1}, {Entity2}, {Entity3}]
-Candidate entities: [{Entity1’}, {Entity2’}, {Entity3’}, …], total {} entities
-Output:
-Table 14. The prompt used to generate entities that are semantically similar to 3 given entities.
-iron, copper, aluminum and zinc.
-math, physics, chemistry and biology.
-{Entity1}, {Entity2}, {Entity3} and ______
-Table 15. The prompt used to generate a class name that covers the given entities.
-Generate a class name that accurately represents the following entities. This class name should encompass all the given entities and reflect their shared characteristics.
-Examples:
-[[[[Tiger, Lion, Cheetah]]]] $\rightarrow$ Big Cats
-[[[[Shakespeare, Tolstoy, Hemingway]]]] $\rightarrow$ Famous Authors
-[[[[Mercury, Venus, Mars]]]] $\rightarrow$ Planets in the Solar System
-[…]
+VII Conclusion
+In conclusion, our study addresses the challenge of ultra-fine-grained Entity Set Expansion (Ultra-ESE) by introducing negative seed entities alongside positive ones, mitigating ambiguity and facilitating the expression of “unwanted” attributes.
+The creation of UltraWiki, a tailored dataset for Ultra-ESE, enables rigorous evaluation and future research.
+Through RetExpan and GenExpan frameworks, we demonstrate the effectiveness of large language models in Ultra-ESE from retrieval-based and generation-based perspectives.
+Our proposed strategies enhance models’ comprehension of ultra-fine-grained entity semantics.
+While our findings are promising, they also highlight the substantial scope for improvement in Ultra-ESE, advancing the field and paving the way for future exploration.
+Acknowledgement
+This research project is supported by National Natural Science Foundation of China (Grant No.62276154), Research Center for ComputerNetwork (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No.2023A1515012914 and 440300241033100801770), Basic Research Fund of Shenzhen City (Grant No.JCYJ20210324120012033, JCYJ20240813112009013 and GJHZ20240218113603006), the Major Key Projectof PCL for Experiments and Applications (PCL2022A05 and PCL2023A09), NSF under Grants III-2106758, and POSE-2346158.
+Xuming Hu was supported by the Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); CAAI-Ant Group Research Fund; Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.